### v1

# Visualized Text-to-Image Retrieval

Abstract
We propose Visualize-then-Retrieve (VisRet), a new paradigm for Text-to-Image (T2I) retrieval that mitigates the limitations of cross-modal similarity alignment of existing multi-modal embeddings. VisRet first projects textual queries into the image modality via T2I generation. Then, it performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features. Experiments on three knowledge-intensive T2I retrieval benchmarks, including a newly introduced multi-entity benchmark, demonstrate that VisRet consistently improves T2I retrieval by 24.5% to 32.7% NDCG@10 across different embedding models. VisRet also significantly benefits downstream visual question answering accuracy when used in retrieval-augmented generation pipelines. The method is plug-and-play and compatible with off-the-shelf retrievers, making it an effective module for knowledge-intensive multi-modal systems. Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
[…]
tcb@breakable
Visualized Text-to-Image Retrieval
Di Wuthanks:  Equal contribution., Yixin Wan11footnotemark: 1, Kai-Wei Chang University of California, Los Angeles {diwu,elaine1wan,kwchang}@cs.ucla.edu
1 Introduction
Figure 1: An overview of VisRet. Compared to the traditional T2I retrieval pipeline, VisRet first projects the text query into the image modality via T2I generation and then performs within-modality retrieval.
Text-to-Image (T2I) retrieval is the task of selecting the most relevant images from a visual corpus based on a textual query. It plays a crucial role in enabling knowledge-intensive applications that require supporting textual inputs with rich visual content (Chen et al., 2022; Wang et al., 2023; Sheynin et al., 2023; Braun et al., 2024).
A common approach to T2I retrieval is to embed both the query and candidate images into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014). However, obtaining accurate similarity rankings that capture fine-grained semantics in both text and image remains a long-standing challenge. Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts”, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023). For instance, Figure 1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer. While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (up-shot). To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025). Yet, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
We propose Visualize-then-Retrieve (VisRet), a novel retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval. Concretely, the textual query is first visualized as one or more images via a T2I generation model. Then, the visualized query, which better exhibits the desired visual-spatial features, is used to perform image-to-image retrieval.
Compared to prior methods, VisRet offers two key advantages. First, visualizations provide a more expressive and intuitive medium for encoding multiple compositional concepts such as entities, poses, and spatial relations, which are difficult to express via text alone. As shown in Figure 1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time. Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in uni-modal retrieval (Koishigarina et al., 2025).
We evaluate VisRet on three challenging T2I retrieval benchmarks: INQUIRE-Rerank (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025), and Visual-RAG-ME, a new benchmark we introduce that features feature comparison questions across multiple entities. Results show that VisRet substantially outperforms baseline T2I retrieval methods and a strong LLM-based query rewriting approach (section 4.2). When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the two baselines by 32.7% and 15.6% higher NDCG@10 respectively, averaged over three benchmarks. With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 24.5% and 12.4%. Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section 4.3). It improves T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval setting and 3.9% and 11.1% in top-10 retrieval setting.
2 Related Work
T2I Retrieval Benchmarks
Early T2I retrieval benchmarks evaluate the ability to identify images based on their paired human-written captions. These datasets span multiple domains and include widely used benchmarks such as Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017). As multi-modal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings. These newer datasets—such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025), and MRAG-Bench (Hu et al., 2024)—shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions. These tasks challenge retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines.
T2I Retrieval Methods
There have been extensive research on improving text-to-image retrieval from different perspectives. First, a number of works aim to train better multi-modal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022). Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025). Finally, a recent line of work introduces generative image retrieval (Li et al., 2024; Qu et al., 2025), which trains a generative model to directly memorize an index of the image corpus. Different from these existing approaches, VisRet expands the query semantics by directly visualizing it in the image space, thereby alleviating the workload of cross-modal retrieval. In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever.
3 Approach
Retrieval Method Visual-RAG Visual-RAG-ME INQUIRE-Rerank-Hard
R@1 R@10 R@30 N@1 N@10 N@30 N@1 N@10 N@30  N@1   N@10 N@30
Retriever = CLIP
Original Query 0.210 0.583 0.737 0.210 0.355 0.385 0.220 0.423 0.435 0.000 0.355 0.412
Query Expansion 0.238 0.586 0.737 0.238 0.360 0.395 0.410 0.575 0.572 0.136 0.349 0.407
Visualize-then-Retrieve 0.251 0.645 0.793 0.251 0.431 0.438 0.460 0.632 0.605 0.170 0.452 0.455
                    - multi-image 0.246 0.637 0.772 0.246 0.414 0.421 0.480 0.629 0.605 0.237 0.428 0.469
Retriever = E5-V
Original Query 0.240 0.568 0.706 0.240 0.386 0.407 0.340 0.465 0.486 0.000 0.319 0.407
Query Expansion 0.223 0.560 0.719 0.223 0.368 0.391 0.460 0.569 0.566 0.170 0.367 0.412
Visualize-then-Retrieve 0.299 0.673 0.801 0.299 0.452 0.461 0.560 0.643 0.622 0.220 0.377 0.425
                   - multi-image 0.307 0.645 0.772 0.307 0.442 0.446 0.520 0.640 0.617 0.203 0.384 0.445
Table 1: Evaluation results across three T2I retrieval benchmarks using different retrieval strategies and retrievers. The best results in each column within each retriever group are boldfaced. R = Recall. N = NDCG.
3.1 Problem Formulation
Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics in $q$. Our paper further considers the task of Visual Question Answering (VQA), where the query is a knowledge-seeking question, with an expected answer $a$.
In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multi-modal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$. Then, a large vision-language model (LVLM) $\mathcal{M}$ directly generates the answer based on the question and the retrieval results $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
3.2 Visualize-then-Retrieve
We introduce Visualize-then-Retrieve (VisRet), a two-staged T2I retrieval pipeline that bridges the modality gap through modality projection. Figure 1 illustrates the pipeline with an intuitive example.
Modality Projection
The first stage of VisRet leverages a T2I generation system $\mathcal{T}$ to directly generate $m$ visualizations $\{v_{1},...,v_{m}\}\equiv\mathcal{T}(q)$. Empirically, we find it helpful to use an LLM within $\mathcal{T}$ to first rephrase $q$ into a T2I instruction $q^{\prime}$, before feeding into existing T2I generation models such as Stable Diffusion (Esser et al., 2024). To generate diverse $\{v_{1},...,v_{m}\}$, randomness can be injected either into $q^{\prime}$ or into the T2I generation.
Within-Modality Retrieval
In the second stage, VisRet performs retrieval within the image modality. Specifically, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},...,r_{k}^{(i)}],$
To aggregate the $m$ separate retrieval results, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009). RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
$\text{score}_{\text{RRF}}(r)=\sum_{i=1}^{m}\frac{1}{\lambda+\text{rank}_{i}(r)},$
where $\text{rank}_{i}(r)$ is the rank position of image $r$ in list $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items. The final top-$k$ retrieval result is formed by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
Dataset Question Ground Truth Baseline Generated Image VisRet
Visual-RAG Does the Mountain Tree Frog (scientific name: Hyla eximia) have any distinctive pattern on the underside of its body? [Uncaptioned image] Rank:49, NDCG@10: 0.00 [Uncaptioned image] Rank:4, NDCG@10: 0.39
How many petals are on each of the Tower Mustard (scientific name: Turritis glabra)’s flowers? [Uncaptioned image] Rank:143, NDCG@10: 0.00 [Uncaptioned image] Rank:2, NDCG@10: 0.76
INQUIRE A male and female cardinal sharing food [Uncaptioned image] Rank:12, NDCG@10: 0.00 [Uncaptioned image] Rank:1, NDCG@10: 1.00
Table 2: Examples: VisRet improves retrieval by highlighting visual features implied by the textual query.
4 Results and Analyses
4.1 Experimental Setup
We evaluate on three challenging benchmarks: (1) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior. We perform additional filtering to remove overly simple queries and call the resulting dataset INQUIRE-Rerank-Hard. (2) Visual-RAG (Wu et al., 2025), a T2I retrieval and VQA benchmark featuring visual knowledge intensive questions on features of natural species that are not commonly documented in text corpus. (3) Visual-RAG-ME, a new benchmark we introduce featuring queries that compare the same visual feature across multiple entities. We present the benchmark details in Appendix A.
For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and NDCG@$k$ with $k$ = 1, 10, 30. For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, following Wu et al. (2025).
For the experiments presented in the main text, we use CLIP and E5-V as the retriever, GPT-4o (OpenAI, 2023b) as the downstream reader, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images. We analyze more model choices in Appendix B and present all the prompts and other hyperparameters in Appendix C.
4.2 Retrieval Performance
Table 1 summarizes retrieval performance across all benchmarks and retrievers. We compare four strategies: using the original textual query, applying query expansion via an LLM, our proposed VisRet, and VisRet with only a single generated image ("-multi-image"). Across all datasets, VisRet consistently outperforms both the original query and query expansion baselines by a large margin. When CLIP is used as the retriever, VisRet outperforms the original query and LLM-based rephrase by 32.7% and 15.6% relatively higher NDCG@10 over three benchmarks. Similar trends hold when E5-V is used as the retriever, exhibiting 24.5% and 12.4% performance gain in NDCG@10. Further, using only one generated image as the query only slightly harms the performance, indicating the flexibility of VisRet. Table 2 presents several examples to demonstrate how the visualization step successfully captures subtle visual semantics implied by the original text query. We present more analyses on the T2I model and the rephrase LLM in Section B.1 and Section B.2.
Figure 2: Downstream RAG-based VQA accuracy on Visual-RAG and Visual-RAG-ME with CLIP as the retriever and GPT-4o as the reader LVLM.
4.3 Downstream QA Performance
To assess the utility of VisRet in real-world applications, we evaluate its downstream VQA accuracy in a RAG pipeline. We compare three settings: (1) using only the model’s internal knowledge, (2) RAG with original text query-based retrieval and (3) RAG with VisRet. Figure 2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever. The original query results in low-quality retrieval augmentation, even slightly harming the performance on Visual-RAG in top-1 retrieval setting compared to no retrieval. By contrast, VisRet significantly improves QA accuracy in both top-1 and top-10 settings on both benchmarks, boosting accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME. Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images containing the required features. Overall, the results confirm that VisRet not only improves retrieval accuracy but also leads to tangible gains in downstream VQA performance. In Section B.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader. In Section B.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
5 Conclusion
This work introduces VisRet, a framework that visualizes text queries to enable more accurate T2I retrieval. By operating entirely in the visual domain during retrieval, VisRet addresses key limitations of cross-modal embedding alignment. Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across three benchmarking datasets and two retrievers. The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multi-modal systems.
Ethics Statement
In this section, we describe the ethical considerations related to this paper.
Potential Risks
Although the goal of this paper is to introduce techniques to improve the text-to-image retrieval performance, the new approach could create new social risks. Specifically, in addition to the embedding model, our approach involves two neural networks: an LLM and a T2I generation model. It is possible for these large models to bring in new social bias in generating the visualized query and thus bias the retrieval results. For instance, when depicting certain scenes of social activity, T2I models could reinforce stereotypical social roles (Wan and Chang, 2024; Wan et al., 2024a). We urge practitioners to implement model debiasing and bias detection measure when deploying our proposed T2I retrieval method in real-world applications. Additionally, previous work has revealed problems with faithfulness and factuality in T2I models (Wan et al., 2024b; tse Huang et al., 2025; Wan and Chang, 2025). We also encourage future studies to explore solutions to similar trustworthy issue with multi-modal generative models.
Artifact Release
Our Visual-RAG-ME annotation is based on Visual-RAG, which is under CC BY-NC 4.0 license and the images shared by the iNaturalist 2021 dataset, which are under one of CC BY 4.0, CC BYNC 4.0, CC BY-NC-ND 4.0, CC BY-NC-SA 4.0, CC0 1.0, CC BY-ND 4.0, CC BY-SA 4.0. We adhere to the intended non-commercial research use of iNaturalist 2021 dataset and do not re-distribute the images. Following Visual-RAG, we will release our Visual-RAG-ME annotations under CC BY-NC 4.0 license.
Acknowledgment
The research is supported in part by Taboola. We also thank Jia-Chen Gu and Hongwei Wang for their valuable feedback.
Supplementary Material: Appendices
Appendix A Benchmark Data Details
In this section, we present the details of Visual-RAG data processing and Visual-RAG-ME annotation. Then, we provide the details of how we processed INQUIRE-Rerenk-Hard. Finally, we document the dataset statistics and the baseline T2I retrieval performance of CLIP (Radford et al., 2021) and E5-V (Jiang et al., 2024) in Table 3.
A.1 Visual-RAG and Visual-RAG-ME
Visual-RAG (Wu et al., 2025) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^1] https://github.com/visual-rag/visual-rag. To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test set. We were able to identify all annotated images in Visual-RAG except a single image due to a likely path error.
We annotate Visual-RAG-ME as an extension of Visual-RAG to visual question answering in the multi-entity setting. Concretely, Visual-RAG-ME reuses the visual features queried in Visual-RAG and constructs questions that compares those features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels. The Visual-RAG-ME annotation pipeline consists of three steps: second entity identification; query composition and filtering; retrieval label annotation and balancing. We next describe the steps in detail.
Second Entity Identification
The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions could be constructed. For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities that have the closest full scientific names.
Query Composition and Filtering
In this step, we (the authors) manually traverse all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question. A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG. The question we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism feature a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.). We were able to construct 82 multi-entity questions after this step. Next, we perform a round of filtering to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions can cover overly similar topics. After the filtering step, we ended up with 50 high quality multi-entity queries.
Retrieval Label Annotation and Balancing
Finally, for each question, we collect images of the second entity from iNaturalist and annotate their retrieval label. A positive label is assigned only if the image clearly displays the feature required to answer the question. For some of the questions, we find that a large number of positive images exist in the iNaturalist database. We therefore implement a filtering step where at most 50 positive images are kept for each entity. Table 8 shows two examples with their questions and ground-truth images.
In Table 3, we present the basic statistics of Visual-RAG-ME. While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
VR VR-ME IR-Hard
# Queries 391 50 59
||||Query|||| (word count) 18.5 25.1 6.0
# Images (per entity) 264 263 100
# Positives (per entity) 14.3 20.8 12.5
CLIP Recall@1 0.210 0.220 0.000
E5-V Recall@1 0.240 0.340 0.000
Table 3: Dataset statistics and baseline performance. VR denotes Visual-RAG and IR denotes INQUIRE-Rerank.
Method R@1 R@10 R@30 N@1 N@10 N@30
Baselines
Original Query 0.210 0.583 0.737 0.210 0.355 0.385
LLM Rephrase 0.238 0.586 0.737 0.238 0.360 0.395
VisRet
DALL-E 3 0.169 0.581 0.757 0.169 0.344 0.376
Stable Diffusion 3 0.166 0.517 0.691 0.166 0.319 0.349
Image-1 (low quality) 0.243 0.611 0.760 0.243 0.397 0.415
Image-1 (high quality) 0.251 0.645 0.793 0.251 0.431 0.438
Table 4: T2I Retrieval performance across different T2I generation models used for VisRet. We use GPT-4o to generate the T2I instruction and CLIP as the retriever. R = Recall. N = NDCG. The best results are boldfaced.
Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
[…]
Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
[…]
Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
Table 5: Retrieval performance on Visual-RAG of with CLIP retriever, using LLMs as T2I instruction generator for VisRet. R = Recall. N = NDCG. The best results are boldfaced.
A.2 INQUIRE-Rerank-Hard
To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (Vendrow et al., 2024) benchmark [^2] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank. The original test set contained 160 queries, each paired with 100 images retrieved by CLIP. In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models. Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1. After manually inspecting the data, we found that for a lot of instances, the negative images are not challenging enough and it is often very straightforward to identify the ground truth images. To highlight the challenging questions, we therefore filtered out the questions on which either CLIP and E5-V can achieve a perfect Recall@1. Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organism themselves, with more challenging confounder negative images.
Knowledge Visual-RAG Visual-RAG-ME
# images GPT-4o-mini GPT-4o GPT-4.1 # images GPT-4o-mini GPT-4o GPT-4.1
Model Knowledge Only 0 38.49 48.47 49.23 0 41.00 51.00 47.00
[…]
10 46.04 51.79 57.06 10 48.00 63.00 65.00
[…]
10 46.42 53.84 56.65 10 55.00 70.00 71.00
Table 6: VQA performance comparison using different LVLMs as instruction generators for VisRet and query rephrase models. CLIP is used as the retriever. Boldfaced numbers indicate the best in each column.
Knowledge Visual-RAG Visual-RAG-ME
# images GPT-4o-mini GPT-4o GPT-4.1 # images GPT-4o-mini GPT-4o GPT-4.1
Model Knowledge Only 0 38.49 48.47 49.23 0 41.00 51.00 47.00
[…]
10 46.42 53.84 56.65 10 55.00 70.00 71.00
Table 7: VQA performance comparison using different knowledge contexts on Visual-RAG and Visual-RAG-ME. CLIP is used as the retriever. Boldfaced numbers indicate the best in each column.
Appendix B VisRet: Further Analyses
In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I generation Model, T2I Instruction LLM, the downstream VQA LVLM reader. Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
B.1 T2I Generation Model Choice
How strong does the T2I generation model need to be for VisRet to work well? We compare the default T2I generation model (Image-1 with high quality setting) with three other models: DALL-E 3 (OpenAI, 2023a), Stable Diffusion 3 (Esser et al., 2024), and Image-1 with the low generation quality setting. Table 4 shows the results with GPT-4o as T2I instruction generation model and CLIP as the retriever model. Interestingly, compared to the cross-modality retrieval baselines, we find that DALL-E 3 and Stable Diffusion 3 do not provide significant performance improvements, while Image-1 low quality clearly and consistently improve the performance. The best performance is achieved by the newest and the most expensive Image-1 high quality setting. Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet. As further T2I generation methods improve, we anticipate that building more cost-efficient version of VisRet is a viable and promising further direction.
B.2 T2I Instruction LLM Choice
Does VisRet work well with other LLMs as the T2I instruction generator? In Table 5, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024). Overall, we observe promising results. For all the three LLMs, using them to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query. While more expensive LLMs achieve a high performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
B.3 Downstream VQA LVLM Choice
While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities? In Table 6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14). Overall, we observe similar trends as those presented in Figure 2. Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former. These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
B.4 Image Queries as Knowledge
As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet. It is a natural question then, that it is still necessary to perform retrieval instead of directly using the generated images as the knowledge? In Table 7, we compare the performance of using a single image as the context with VisRet. Overall, we observe a mixed result. For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern. For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefers the generated image over top-1 retrieval (and even top-10 retrieval). However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than using the generated image. Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-MR and cannot be fully replaced by pure T2I generation at this stage. It is an important future work to further combine image generation and image retrieval to improve the quality of the retrieved knowledge.
B.5 Further Qualitative Studies
In Table 8, we additionally provide some qualitative examples of how VisRet successfully improves T2I retrieval performance. As can be observed in top-ranked examples in text query-only retrieval, these retrieved images often fall short in correctly conveying visual semantics information such as angle, body part depicted, form of the subject depicted, visual distance of the subject, and so on. In contrast, VisRet is able to represent the inferred nuanced visual semantics in the image generation step first, then utilize this within-modality semantic-rich image query to obtain precise visual knowledge related to the VQA task.
Appendix C Implementation Details
In this section, we present the implementation details of VisRet and baselines.
VisRet: T2I Generation
To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in it. The prompts for three benchmarking datasets are shown in Figure 3, Figure 4, and Figure 5, respectively. Then, we wrap the rephrased query with a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation. We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high". For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a signle instruction already results in images generated with a high level of diversity. Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
VisRet: Retrieval
After obtaining the generated visualizations, we encode both the visualized images and the image corpus via an off-the-shelf CLIP [^3] https://huggingface.co/openai/clip-vit-large-patch14-336 or E5-V [^4] https://huggingface.co/royokong/e5-v encoder and perform a similarity search. Cosine similarity is used as the similarity metric. For RRF, we used $\lambda=1$ to merge the rankings from multiple queries. All the retrieval experiments were performed on a local server with Nvidia A100 GPU.
VQA Answer Generation
We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (Wei et al., 2022). Concretely, the model is asked to always extract visual information, perform reasoning, and conclude its reasoning with self-verification. We show the detailed prompt in Figure 6 and Figure 7.
Baselines
For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important feature that the query is seeking. At the early stage of the prompt, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction. Therefore, we report the results with the same prompt for the final version for the paper.
VQA Evaluation
For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025), as shown in Figure 8. Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering. We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
Dataset Question Ground Truth Baseline Generated Image VisRet
Visual-RAG What is the color of the head of larva Urania Swallowtail Moth (scientific name: Urania fulgens)? [Uncaptioned image] Rank:96, NDCG@10: 0.00 [Uncaptioned image] Rank:1, NDCG@10: 0.66
What color are the ventral abdomen of Golden Buprestid Beetle (scientific name: Buprestis aurulenta)? [Uncaptioned image] Rank:23, NDCG@10: 0.00 [Uncaptioned image] Rank:4, NDCG@10: 0.39
Are there visually distinctive scales on feet of Azure Tit (scientific name: Cyanistes cyanus)? [Uncaptioned image] Rank:26, NDCG@10: 0.40 [Uncaptioned image] Rank:2, NDCG@10: 0.76
INQUIRE A male and female cardinal sharing food [Uncaptioned image] Rank:12, NDCG@10: 0.00 [Uncaptioned image] Rank:1, NDCG@10: 1.00
Mexican grass-carrying wasp visiting a purple flower [Uncaptioned image] Rank:45, NDCG@10: 0.00 [Uncaptioned image] Rank:4, NDCG@10: 0.39
great golden digger wasp carrying an orthopteron [Uncaptioned image] Rank:12, NDCG@10: 0.45 [Uncaptioned image] Rank:2, NDCG@10: 0.83
male ruby-throated hummingbird in flight [Uncaptioned image] Rank:60, NDCG@10: 0.00 [Uncaptioned image] Rank:7, NDCG@10: 0.39
Visual-RAG-ME Which one has striped primary flight feathers, Willet (scientific name: Tringa semipalmata) or Grey-tailed tattler (scientific name: Tringa brevipes)? [Uncaptioned image] Entity: Willet Rank:104, NDCG@10: 0.00 [Uncaptioned image] Entity: Willet Rank:1, NDCG@10: 0.72
[Uncaptioned image] Entity: Grey-tailed tattler Rank:74, NDCG@10: 0.00 [Uncaptioned image] Entity: Grey-tailed tattler Rank:1, NDCG@10: 1.00
Which one has less prominent color patterns, silvery checkerspot (scientific name: Chlosyne nycteis) caterpillar or theona checkerspot (scientific name: Chlosyne theona) caterpillar? [Uncaptioned image] Entity: silvery checkerspot Rank:188, NDCG@10: 0.45 [Uncaptioned image] Entity: silvery checkerspot Rank:4, NDCG@10: 0.88
[Uncaptioned image] Entity: theona checkerspot Rank:120, NDCG@10: 0.52 [Uncaptioned image] Entity: theona checkerspot Rank:3, NDCG@10: 0.85
Table 8: Additional qualitative results on the three benchmarking datasets.
You are given a query, rephrase the query into a short descriptive phrase that highlights the key part of the entity where the queried feature could be found. DO NOT include the asked feature (shape, color, etc.) but instead include the part of the entity where the feature could be found. Output only the rephrased query. Examples: Original query: What shape are the flowers of bush flax (scientific name: Astelia fragrans)? Rephrased query: flowers of bush flax Original query: Is the any specific color pattern on underside wings of tawny pipit (scientific name: Anthus campestris) displayed during flight, or is it uniformly colored? Rephrased query: tawny pipit with its underside wings shown Original query: {question} Rephrased query:
Figure 3: Prompt for instructing an LLM to generate the T2I generation instruction for Visual-RAG questions.
You are given a query about two entities, as well as an entity of interest. Rephrase the query into a short descriptive phrase that highlights the key part of the entity of interest on which the queried feature could be found. DO NOT include the asked feature (shape, color, etc.) but instead include the entity name + part of the entity where the feature could be found. Output only the rephrased query. Examples: Original query: Are the tongues of grass snake (scientific name: Natrix helvetica) and Chicken Snake (scientific name: Spilotes pullatus) the same color? Entity of interest: Spilotes pullatus Rephrased query: Chicken Snake with its tongue shown Original query: Which one has a more slender matured legume, common milkpea (scientific name: Galega officinalis) or narrowleaf lupin (scientific name: Lupinus angustifolius)? Entity of interest: Galega officinalis Rephrased query: the legume of common milkpea Original query: {question} Entity of interest: {entity} Rephrased query:
Figure 4: Prompt for instructing an LLM to generate the T2I generation instruction for Visual-RAG-ME questions.
You are given an image retrieval query, rephrase the query to add in a bit detail (no longer than 30 words). The rephrased query should highlight the appearance, posture, actions of the main entity so that it is easier to retrieve the desired image among (1) images of the same entity with different posture and (2) images of different entities with the same posture. Original query: {question} Rephrased query:
Figure 5: Prompt for instructing an LLM to generate the T2I generation instruction for INQUIRE-Rerank questions.
[…]
Figure 6: Prompt for VQA on Visual-RAG.
[…]
Figure 7: Prompt for VQA on Visual-RAG-ME.
Please evaluate the answer to a question, score from 0 to 1. The reference answer is provided, and the reference is usually short phrases or a single keyword. If the student answer is containing the keywords or similar expressions (including similar color), without any additional guessed information, it is full correct. If the student answer have missed some important part in the reference answer, please assign partial score. Usually, when there are 2 key features and only 1 is being answered, assign 0.5 score; if there are more than 2 key features, adjust partial score by ratio of correctly answered key feature. The reference answer can be in the form of a Python list, in this case, any one of the list item is correct. If student answer contain irrelevant information not related to question, mark it with "Redundant", but it does not affect score if related part are correct. (e.g. Question: what shape is leave of Sanguinaria canadensis, Student Answer: shape is xxx, color is yyy, this is Redundant answer) If student answer contain features not listed in reference answer, mark it with "Likely Hallucination" and deduct 0.5 score. (e.g., Reference Answer: black and white. Student Answer: black white, with yellow dots, “yellow dots” is not mentioned in reference). The reference answer sometimes contains an add-on enclosed by brackets (), to help verifying hallucinations (e.g.: “shape is xxx (color is yyy)”). Not mentioning add-on information in answer is not considered wrong. Answering "I don’t know", "Not enough information" is considered wrong. Format Instructions: Separate the remarks with score using "|", that is, use the syntax of: "Score: {score} | Likely Hallucination", "Score: {score}", "Score: {score} | Likely Hallucination | Redundant". If any explanation on why giving the score is needed, do not start a new line and append after remark with brackets, e.g. "Score: {score} | Redundant | (Explanation: abc)". Following are few examples: Question: Is there any specific color marking around the eyes of a semipalmated plover (scientific name: Charadrius semipalmatus)? Reference Answer: black eye-round feather, white stripe above eyes. (sometimes connected to the white forehead) Student Answer: Yes, the bird has a distinctive black line that runs through the eye, which is a key identifying feature. Score: 0 | Likely Hallucination Student Answer: They have a black vertical band in front of the eye, a white band above the eye, and a single black band that wraps partially around the eye, creating a partial "mask" appearance. Score: 1 Student Answer: Yes, the semipalmated plover has a distinctive black/dark ring around its eye, surrounded by a bright white ring or patch Score: 0.5 | Likely Hallucination (Explanation: not white ring, but only a line above the eye) Question: What is the typical color of the antennae of Harris’s checkerspot butterfly (scientific name: Chlosyne harrisii)? Reference Answer: alternating black and white band, with yellow on the tip Student Answer: The antennae of Harris’s checkerspot butterfly are black with orange-tipped clubs. Score: 0.5 (Explanation: not mentioning black and white) Student Answer: The typical color of the antennae of Harris’s checkerspot butterfly is black with white spots. Score: 0.5 | Likely Hallucination (Explanation: not white spot but band. Not mentioning the tip) Question: Are the leaves of burro-weed (scientific name: Ambrosia dumosa) usually covered in small hairs? Reference Answer: yes Student Answer: Yes, the leaves of burro-weed (Ambrosia dumosa) are typically covered in small hairs, giving them a grayish or whitish-green appearance. Score: 1 | Redundant Now, score the following question: Question: {question} Reference Answer: {reference_answer} Student Answer: {model_answer}
Figure 8: Prompt for the LLM VQA judge used for Visual-RAG and Visual-RAG-ME.

### v2

@@
-Visualized Text-to-Image Retrieval
+VisRet: Visualization Improves Knowledge-Intensive Text-to-Image Retrieval
 Abstract
-We propose Visualize-then-Retrieve (VisRet), a new paradigm for Text-to-Image (T2I) retrieval that mitigates the limitations of cross-modal similarity alignment of existing multi-modal embeddings.
+Text-to-image retrieval (T2I retrieval) remains challenging because cross-modal embeddings often behave as bags of concepts and underrepresent structured visual relationships such as pose and viewpoint.
+We propose Visualize-then-Retrieve (VisRet), a new paradigm for T2I retrieval that mitigates this limitation of cross-modal similarity alignment.
 VisRet first projects textual queries into the image modality via T2I generation.
 Then, it performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features.
-Experiments on three knowledge-intensive T2I retrieval benchmarks, including a newly introduced multi-entity benchmark, demonstrate that VisRet consistently improves T2I retrieval by 24.5% to 32.7% NDCG@10 across different embedding models.
-VisRet also significantly benefits downstream visual question answering accuracy when used in retrieval-augmented generation pipelines.
-The method is plug-and-play and compatible with off-the-shelf retrievers, making it an effective module for knowledge-intensive multi-modal systems.
+Across four benchmarks (Visual-RAG, INQUIRE-Rerank, Microsoft COCO, and our new Visual-RAG-ME featuring multi-entity comparisons), VisRet substantially outperforms cross-modal similarity matching and baselines that recast T2I retrieval as text-to-text similarity matching, improving nDCG@30 by 0.125 on average with CLIP as the retriever and by 0.121 with E5-V.
+For downstream question answering, VisRet increases accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval, and by 3.9% and 11.1% in top-10 retrieval.
+Ablation studies show compatibility with different T2I instruction LLMs, T2I generation models, and downstream LLMs.
+VisRet provides a practical and principled path that energizes further advances in vision-language retrieval.
 Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
-[…]
-tcb@breakable
-Visualized Text-to-Image Retrieval
-Di Wuthanks:  Equal contribution., Yixin Wan11footnotemark: 1, Kai-Wei Chang University of California, Los Angeles {diwu,elaine1wan,kwchang}@cs.ucla.edu
+VisRet: Visualization Improves Knowledge-Intensive
+Text-to-Image Retrieval
+Yixin Wan1thanks:  Equal contribution., Di Wu [^1] footnotemark: 1, Kai-Wei Chang1 1University of California, Los Angeles {diwu,kwchang}@cs.ucla.edu
 1 Introduction
+Text-to-Image (T2I) retrieval is the task of selecting the most relevant images from a visual corpus based on a textual query.
+It plays a crucial role in enabling knowledge-intensive applications that require supporting textual inputs with rich visual content (Chen et al., 2022; Wang et al., 2023; Sheynin et al., 2023; Braun et al., 2024).
+A common approach to T2I retrieval is to embed both the text query and image candidates into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014).
+However, obtaining accurate similarity rankings that capture fine-grained semantic alignments between text and image remains a long-standing challenge.
+Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts” for similarity matching, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023).
+For instance, Figure˜1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer.
+While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (upward-facing shot).
+To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025).
+Nevertheless, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
 Figure 1: An overview of VisRet.
 Compared to the traditional T2I retrieval pipeline, VisRet first projects the text query into the image modality via T2I generation and then performs within-modality retrieval.
-Text-to-Image (T2I) retrieval is the task of selecting the most relevant images from a visual corpus based on a textual query.
-It plays a crucial role in enabling knowledge-intensive applications that require supporting textual inputs with rich visual content (Chen et al., 2022; Wang et al., 2023; Sheynin et al., 2023; Braun et al., 2024).
-A common approach to T2I retrieval is to embed both the query and candidate images into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014).
-However, obtaining accurate similarity rankings that capture fine-grained semantics in both text and image remains a long-standing challenge.
-Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts”, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023).
-For instance, Figure 1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer.
-While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (up-shot).
-To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025).
-Yet, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
 We propose Visualize-then-Retrieve (VisRet), a novel retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval.
-Concretely, the textual query is first visualized as one or more images via a T2I generation model.
-Then, the visualized query, which better exhibits the desired visual-spatial features, is used to perform image-to-image retrieval.
+Concretely, VisRet first utilizes a text-to-image generation (T2I generation) model to visualize the textual query, highlighting the key desired visual-spatial features.
+Then, the visualized query is used to perform direct image-to-image retrieval.
 Compared to prior methods, VisRet offers two key advantages.
-First, visualizations provide a more expressive and intuitive medium for encoding multiple compositional concepts such as entities, poses, and spatial relations, which are difficult to express via text alone.
-As shown in Figure 1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time.
-Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in uni-modal retrieval (Koishigarina et al., 2025).
-We evaluate VisRet on three challenging T2I retrieval benchmarks: INQUIRE-Rerank (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025), and Visual-RAG-ME, a new benchmark we introduce that features feature comparison questions across multiple entities.
-Results show that VisRet substantially outperforms baseline T2I retrieval methods and a strong LLM-based query rewriting approach (section 4.2).
-When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the two baselines by 32.7% and 15.6% higher NDCG@10 respectively, averaged over three benchmarks.
-With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 24.5% and 12.4%.
-Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section 4.3).
+First, visualizations provide a more expressive and intuitive medium for representing multiple compositional concepts, including entities, poses, and spatial relations.
+These concepts are often difficult to convey through text alone, and attempting to encode them exhaustively in a text query can impair cross-modal matching performance due to limitations in embedding quality.
+In contrast, as shown in Figure˜1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time.
+Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in unimodal retrieval (Koishigarina et al., 2025).
+To investigate the effectiveness of VisRet, we conduct exhaustive experiments on two commonly used cross-modal embedding models using four benchmarks: Visual-RAG (Wu et al., 2025b), INQUIRE-Rerank (Vendrow et al., 2024), Microsoft COCO (Chen et al., 2015), and Visual-RAG-ME, a new benchmark we introduce that features challenging visual feature comparison questions across multiple entities.
+Results show that VisRet substantially outperforms baseline cross-modal similarity matching as well as three strong baseline methods that casts T2I retrieval to text-to-text similarity matching (section˜5.2).
+When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the baselines by 0.125 higher nDCG@30, averaged over four benchmarks.
+With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 0.121.
+Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section˜5.3).
 It improves T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval setting and 3.9% and 11.1% in top-10 retrieval setting.
+Further analyses show that VisRet is compatible with different T2I instruction generation and downstream question answering LLMs.
+In addition, the retrieval performance consistently improves with stronger T2I models.
+Finally, although they are effective in retrieving visual knowledge, the generated image queries themselves are generally insufficient as standalone knowledge to answer downstream questions.
 2 Related Work
@@
 Early T2I retrieval benchmarks evaluate the ability to identify images based on their paired human-written captions.
-These datasets span multiple domains and include widely used benchmarks such as Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017).
-As multi-modal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings.
-These newer datasets—such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025), and MRAG-Bench (Hu et al., 2024)—shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions.
-These tasks challenge retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines.
+These datasets span multiple domains and include widely used benchmarks such as COCO (Chen et al., 2015), Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017).
+As multimodal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings.
+These newer datasets, such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025b), and MRAG-Bench (Hu et al., 2024), shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions.
+These tasks challenge T2I retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines.
+Further extending this stream of work, our Visual-RAG-ME benchmarks T2I retrieval for reasoning over the visual features across multiple entities.
 T2I Retrieval Methods
-There have been extensive research on improving text-to-image retrieval from different perspectives.
-First, a number of works aim to train better multi-modal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022).
-Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025).
+Extensive research has focused on improving text-to-image retrieval from different perspectives.
+First, a number of works aim to train better multimodal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022).
+Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025; Ding et al., 2025).
 Finally, a recent line of work introduces generative image retrieval (Li et al., 2024; Qu et al., 2025), which trains a generative model to directly memorize an index of the image corpus.
 Different from these existing approaches, VisRet expands the query semantics by directly visualizing it in the image space, thereby alleviating the workload of cross-modal retrieval.
-In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever.
+In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever and pre-computed image embedding indices.
 3 Approach
-Retrieval Method Visual-RAG Visual-RAG-ME INQUIRE-Rerank-Hard
-R@1 R@10 R@30 N@1 N@10 N@30 N@1 N@10 N@30  N@1   N@10 N@30
-Retriever = CLIP
-Original Query 0.210 0.583 0.737 0.210 0.355 0.385 0.220 0.423 0.435 0.000 0.355 0.412
-Query Expansion 0.238 0.586 0.737 0.238 0.360 0.395 0.410 0.575 0.572 0.136 0.349 0.407
-Visualize-then-Retrieve 0.251 0.645 0.793 0.251 0.431 0.438 0.460 0.632 0.605 0.170 0.452 0.455
-                    - multi-image 0.246 0.637 0.772 0.246 0.414 0.421 0.480 0.629 0.605 0.237 0.428 0.469
-Retriever = E5-V
-Original Query 0.240 0.568 0.706 0.240 0.386 0.407 0.340 0.465 0.486 0.000 0.319 0.407
-Query Expansion 0.223 0.560 0.719 0.223 0.368 0.391 0.460 0.569 0.566 0.170 0.367 0.412
-Visualize-then-Retrieve 0.299 0.673 0.801 0.299 0.452 0.461 0.560 0.643 0.622 0.220 0.377 0.425
-                   - multi-image 0.307 0.645 0.772 0.307 0.442 0.446 0.520 0.640 0.617 0.203 0.384 0.445
-Table 1: Evaluation results across three T2I retrieval benchmarks using different retrieval strategies and retrievers.
-The best results in each column within each retriever group are boldfaced.
-R = Recall.
-N = NDCG.
 3.1 Problem Formulation
-Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics in $q$.
+Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics of $q$.
 Our paper further considers the task of Visual Question Answering (VQA), where the query is a knowledge-seeking question, with an expected answer $a$.
-In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multi-modal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$.
+In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multimodal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$.
 Then, a large vision-language model (LVLM) $\mathcal{M}$ directly generates the answer based on the question and the retrieval results $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
 3.2 Visualize-then-Retrieve
-We introduce Visualize-then-Retrieve (VisRet), a two-staged T2I retrieval pipeline that bridges the modality gap through modality projection.
-Figure 1 illustrates the pipeline with an intuitive example.
+We introduce Visualize-then-Retrieve (VisRet), a two-stage T2I retrieval pipeline that bridges the modality gap through modality projection.
+Figure˜1 illustrates the pipeline with an intuitive example.
 Modality Projection
 The first stage of VisRet leverages a T2I generation system $\mathcal{T}$ to directly generate $m$ visualizations $\{v_{1},...,v_{m}\}\equiv\mathcal{T}(q)$.
-Empirically, we find it helpful to use an LLM within $\mathcal{T}$ to first rephrase $q$ into a T2I instruction $q^{\prime}$, before feeding into existing T2I generation models such as Stable Diffusion (Esser et al., 2024).
-To generate diverse $\{v_{1},...,v_{m}\}$, randomness can be injected either into $q^{\prime}$ or into the T2I generation.
+Concretely, based on the original query, an LLM first operates in the text space to draft T2I instruction $q^{\prime}$, describing the images that likely satisfy the feature requirements implied in $q$.
+Next, we use an existing T2I generation method, such as Stable Diffusion (Esser et al., 2024), to project $q^{\prime}$ into a list of image queries $\{v_{1},...,v_{m}\},v_{i}\in\mathcal{I}$.
+To generate diverse images, randomness can be injected either into $q^{\prime}$ or into the T2I generation.
+In this work, we sample $m$ times with the same $q^{\prime}$.
 Within-Modality Retrieval
-In the second stage, VisRet performs retrieval within the image modality.
-Specifically, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
+In the second stage, VisRet performs retrieval exclusively within the image modality.
+First, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
 $\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},...,r_{k}^{(i)}],$
-To aggregate the $m$ separate retrieval results, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009).
-RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
+Then, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) to aggregate the $m$ retrieval results. RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
 $\text{score}_{\text{RRF}}(r)=\sum_{i=1}^{m}\frac{1}{\lambda+\text{rank}_{i}(r)},$
 where $\text{rank}_{i}(r)$ is the rank position of image $r$ in list $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items.
-The final top-$k$ retrieval result is formed by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
-Dataset Question Ground Truth Baseline Generated Image VisRet
+The final top-$k$ retrieval result is obtained by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
+4 Visual-RAG-ME Benchmark
+To facilitate the design of intelligent T2I retrieval systems, we introduce a multi-entity question answering and T2I retrieval dataset Visual-RAG-ME as an extension of Visual-RAG.
+Concretely, Visual-RAG-ME constructs questions that compare the visual features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels.
+The Visual-RAG-ME annotation pipeline consists of three steps as described below.
+Second Entity Identification
+The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions can be constructed.
+For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities with the most similar scientific names.
+Query Composition and Filtering
+In this step, we (the authors) manually review all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question.
+A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG.
+The questions we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism features a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.).
+We were able to construct 82 multi-entity questions after this step.
+Then, a round of filtering is performed to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions that cover overly similar topics.
+The final Visual-RAG-ME contains 50 high-quality multi-entity queries.
+Retrieval Label Annotation and Balancing
+For each question, we collect images of the second entity from the iNaturalist 2021 dataset (Horn et al., 2021) and annotate their retrieval labels.
+A positive label is assigned only if the image clearly displays the feature required to answer the question.
+For some of the questions, we find that a large number of positive images exist in the iNaturalist database.
+We therefore implement a filtering step where at most 50 positive images are kept for each entity.
+Table˜3 and Appendix Table˜8 show examples with their questions and ground-truth images.
+In Table˜1, we present the basic statistics of Visual-RAG-ME.
+While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
+VR VR-ME IR-Hard $C$-Hard
+# Queries 391 50 59 500
+||Query|| (word count) 18.5 25.1 6.0 12.2
+# Images (per entity) 264 263 100 5000
+# Positives (per entity) 14.3 20.8 12.5 1
+CLIP Recall@1 0.210 0.220 0.000 0.000
+E5-V Recall@1 0.240 0.340 0.000 0.000
+Table 1: Dataset statistics and baseline performance. VR = Visual-RAG, IR = INQUIRE-Rerank, and $C$ = Microsoft COCO 2017.
+Retrieval Method Visual-RAG Visual-RAG-ME INQUIRE-Rerank-Hard COCO-Hard
+N@1 N@10 N@30 N@1 N@10 N@30 N@1 N@10 N@30 N@1 N@10 N@30
+Retriever = CLIP
+Original Query 0.210 0.355 0.385 0.220 0.423 0.435 0.000 0.355 0.412 0.000 0.017 0.042
+LLM Rewriting 0.238 0.360 0.395 0.410 0.575 0.572 0.136 0.349 0.407 0.014 0.047 0.093
+Corpus Captioning (BLIP) 0.105 0.217 0.271 0.210 0.354 0.371 0.136 0.319 0.401 0.054 0.122 0.153
+VISA Reranking (top-30) 0.223 0.378 0.388 0.240 0.443 0.457 0.000 0.000 0.000 0.000 0.000 0.000
+Captioning Reranking (top-30) 0.200 0.336 0.367 0.160 0.362 0.390 0.000 0.000 0.000 0.000 0.000 0.000
+VisRet 0.251 0.431 0.438 0.460 0.632 0.605 0.170 0.452 0.455 0.034 0.072 0.108
+VisRet - multi-image 0.246 0.414 0.421 0.480 0.629 0.605 0.237 0.428 0.469 0.038 0.069 0.106
+Retriever = E5-V
+Original Query 0.240 0.386 0.407 0.340 0.465 0.486 0.000 0.319 0.407 0.000 0.163 0.178
+LLM Rewriting 0.223 0.368 0.391 0.460 0.569 0.566 0.170 0.367 0.412 0.038 0.133 0.182
+Corpus Captioning (BLIP) 0.156 0.294 0.319 0.240 0.405 0.436 0.119 0.342 0.398 0.064 0.158 0.196
+VISA Reranking (top-30) 0.279 0.413 0.419 0.350 0.495 0.500 0.022 0.002 0.002 0.000 0.000 0.000
+Captioning Reranking (top-30) 0.248 0.369 0.396 0.400 0.429 0.454 0.000 0.000 0.005 0.000 0.000 0.000
+VisRet 0.299 0.452 0.461 0.560 0.643 0.622 0.220 0.377 0.425 0.042 0.173 0.205
+VisRet - multi-image 0.307 0.442 0.446 0.520 0.640 0.617 0.203 0.384 0.445 0.048 0.168 0.206
+Table 2: nDCG (N@k) across four T2I retrieval benchmarks using different retrieval strategies and retrievers.
+Within each retriever group, the best results per column are bolded.
+We use different colors to denote similarity matching in text-to-image, text-to-text, or image-to-image style.
+VisRet achieves the state-of-the-art across most evaluation settings, especially excelling in handling knowledge-intensive queries from Visual-RAG and Visual-RAG-ME.
+Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
 Visual-RAG Does the Mountain Tree Frog (scientific name: Hyla eximia) have any distinctive pattern on the underside of its body?
@@
 How many petals are on each of the Tower Mustard (scientific name: Turritis glabra)’s flowers? [Uncaptioned image] Rank:143, nDCG@10: 0.00 [Uncaptioned image] Rank:2, nDCG@10: 0.76
+Visual-RAG-ME Which one has striped primary flight feathers, Willet (scientific name: Tringa semipalmata) or Grey-tailed tattler (scientific name: Tringa brevipes)?
+[Uncaptioned image] Entity: Willet Rank:104, nDCG@10: 0.00 [Uncaptioned image] Entity: Willet Rank:1, nDCG@10: 0.72
+[Uncaptioned image] Entity: Grey-tailed tattler Rank:74, nDCG@10: 0.00 [Uncaptioned image] Entity: Grey-tailed tattler Rank:1, nDCG@10: 1.00
 INQUIRE A male and female cardinal sharing food [Uncaptioned image] Rank:12, nDCG@10: 0.00 [Uncaptioned image] Rank:1, nDCG@10: 1.00
+COCO Plate with pizza, knife and fork laying on edge of plate with two glasses next to them. [Uncaptioned image] Rank:39, nDCG@10: 0.00 [Uncaptioned image] Rank:3, nDCG@10: 0.3937
 Table 3: Examples: VisRet improves retrieval by highlighting visual features implied by the textual query.
-4 Results and Analyses
+"Baseline" refers to cross-modal retrieval with the original query.
+CLIP embedding scores are used for this visualization.
+5 Main Results
 5.1 Experimental Setup
-We evaluate on three challenging benchmarks: (1) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior.
-We perform additional filtering to remove overly simple queries and call the resulting dataset INQUIRE-Rerank-Hard.
-(2) Visual-RAG (Wu et al., 2025), a T2I retrieval and VQA benchmark featuring visual knowledge intensive questions on features of natural species that are not commonly documented in text corpus.
+Benchmarks
+We evaluate on four challenging benchmarks: (1) Visual-RAG (Wu et al., 2025b), a T2I retrieval and VQA benchmark featuring visual knowledgeintensive questions on features of natural species that are not commonly documented in text corpora.
 (2) Visual-RAG-ME, a new benchmark we introduce featuring queries that compare the same visual feature across multiple entities.
-We present the benchmark details in Appendix A.
-For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and NDCG@$k$ with $k$ = 1, 10, 30.
-For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, following Wu et al.
-(2025).
-For the experiments presented in the main text, we use CLIP and E5-V as the retriever, GPT-4o (OpenAI, 2023b) as the downstream reader, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images.
-We analyze more model choices in Appendix B and present all the prompts and other hyperparameters in Appendix C.
-4.2 Retrieval Performance
-Table 1 summarizes retrieval performance across all benchmarks and retrievers.
-We compare four strategies: using the original textual query, applying query expansion via an LLM, our proposed VisRet, and VisRet with only a single generated image ("-multi-image").
-Across all datasets, VisRet consistently outperforms both the original query and query expansion baselines by a large margin.
-When CLIP is used as the retriever, VisRet outperforms the original query and LLM-based rephrase by 32.7% and 15.6% relatively higher NDCG@10 over three benchmarks.
-Similar trends hold when E5-V is used as the retriever, exhibiting 24.5% and 12.4% performance gain in NDCG@10.
-Further, using only one generated image as the query only slightly harms the performance, indicating the flexibility of VisRet.
-Table 2 presents several examples to demonstrate how the visualization step successfully captures subtle visual semantics implied by the original text query.
-We present more analyses on the T2I model and the rephrase LLM in Section B.1 and Section B.2.
+(3) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior.
+(4) Microsoft COCO 2017 (Chen et al., 2015), a widely used image captioning dataset, which we repurpose for caption-to-image retrieval evaluation.
+For (3) and (4), we perform additional filtering to retain the most difficult queries and call the resulting datasets INQUIRE-Rerank-Hard and COCO-Hard.
+Table˜1 shows the basic statistics of each dataset.
+We present additional details in Appendix Appendix˜A.
+Metrics
+For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and nDCG@$k$ with $k$ = 1, 10, 30111As nDCG@1 is undefined, we use Recall@1 as its value..
+For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, using a prompt derived from Visual-RAG (Wu et al., 2025b).
+Baselines
+Using the benchmarks, we evaluate three families of T2I retrieval methods:
+  1. 1.
+    Text-to-Image Matching: Retrieving with the original textual query against the image index using off-the-shelf retrievers (CLIP, E5-V).
+We also compare with a baseline of applying query rewriting via an LLM to highlight the desired features and then performing standard T2I retrieval over the same image embedding index.
+  2. 2.
+    Text-to-Text Matching: Converting the images into textual surrogates for text-to-text retrieval.
+We compare with (1) captioning the entire corpus with BLIP, (2) top-k reranking with VISA, and (3) an ablated version of VISA, where we rerank top-k images using LLM-generated captions.
+We use Qwen2.5-70B-VL (Bai et al., 2025) as the model for (2) and (3) and rerank the top-30 retrieval.
+  3. 3.
+    Image-to-Image Matching: We evaluate VisRet and an image ablation of VisRet that only uses a single image instead of multiple images (-multi-image).
+Implementation Details
+For all the retrieval experiments, we use CLIP and E5-V as the retrievers, GPT-4o (OpenAI, 2023b) as T2I instruction generation LLM, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images.
+For the downstream question answering (Visual-RAG and Visual-RAG-ME), we use GPT-4o as the reader LLM.
+We present the full implementation details of VisRet and the baselines in Appendix˜B.
+5.2 VisRet Improves T2I Retrieval
+[…]
+Compared with VisRet, the text-to-text retrieval baselines exhibit inconsistent performance across the benchmarks.
+Captioning the entire corpus with BLIP only works on COCO, which has the closest relationship to its training data, while it does not improve performance over cross-modal retrieval on the other benchmarks with more knowledge intensive queries.
+We hypothesize that this is caused by the high information density in images, which incurs an inevitable loss of potentially useful information during captioning.
+On the other hand, with access to cross-modal retrieval results and by directly operating on them, Captioning Reranking and VISA outperform query-agnostic corpus captioning, yet cannot outperform direct cross-modal retrieval by a large margin.
+Furthermore, these methods can only process a limited number of top-k retrieved items as their cost grows linearly with k.
+As a result, their performance is limited by the first round of retrieval, leading to near-zero performance on INQUIRE-Rerank-Hard and COCO-Hard.
+Finally, comparing VisRet with its single-image variant, we find that using only one generated image as the query slightly harms the performance but still outperforms most of the other baselines.
+The finding holds for both the lightweight CLIP embedding or the stronger E5-V embeddings, indicating the flexibility of VisRet.
+Table˜3 presents several examples to demonstrate how the visualization step successfully captures subtle visual features implied by the original text query.
+As shown in the examples, for a number of ground truth images, the main entity is either in a particular posture that reveals the required information (Visual-RAG and Visual-RAG-ME) or in a complex relationships between other entities (INQUIRE and COCO), posing a challenge to cross-modal similarity matching to recognize and satisfy these constraints.
+On the other hand, VisRet infuses these constraints naturally into the image query and significantly reduces the workload of the downstream embedding model.
 Figure 2: Downstream RAG-based VQA accuracy on Visual-RAG and Visual-RAG-ME with CLIP as the retriever and GPT-4o as the reader LVLM.
-4.3 Downstream QA Performance
+5.3 VisRet improves Downstream QA
+Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
+Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
+LLM Rewriting Llama 3.1 8B Instruct 0.238 0.563 0.737 0.238 0.365 0.385
+Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
+GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
+Visualize-then-Retrieve Llama 3.1 8B Instruct 0.243 0.606 0.780 0.243 0.405 0.428
+Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
+GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
+Table 4: Retrieval performance on Visual-RAG of with CLIP retriever, using different LLMs as T2I instruction generator for VisRet. R = Recall. N = nDCG. The best results are boldfaced.
 To assess the utility of VisRet in real-world applications, we evaluate its downstream VQA accuracy in a RAG pipeline.
 We compare three settings: (1) using only the model’s internal knowledge, (2) RAG with original text query-based retrieval and (3) RAG with VisRet.
-Figure 2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever.
-The original query results in low-quality retrieval augmentation, even slightly harming the performance on Visual-RAG in top-1 retrieval setting compared to no retrieval.
+Figure˜2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever.
+The original query results in low-quality retrieval augmentation, even slightly harming performance on Visual-RAG in top-1 retrieval setting compared to no retrieval.
 By contrast, VisRet significantly improves QA accuracy in both top-1 and top-10 settings on both benchmarks, boosting accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME.
-Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images containing the required features.
+Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images that contain the required features.
 Overall, the results confirm that VisRet not only improves retrieval accuracy but also leads to tangible gains in downstream VQA performance.
-In Section B.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader.
-In Section B.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
+In Section˜6.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader.
+In Section˜6.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
+6 Analyses
+In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I instruction LLM, T2I generation model, and the downstream VQA reader.
+Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
+6.1 Query Generation Model Choice
+Does VisRet work well with other LLMs as the T2I instruction generator?
+In Table˜4, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024).
+Overall, we observe promising results: For all three LLMs, using the LLMs themselves to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query.
+While more expensive LLMs achieve higher performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
+Method N@1 N@10 N@30 Avg
+Retriever = CLIP
+Original Query 0.220 0.423 0.435 0.359
+VisRet Variants:
+Stable Diffusion 3.5 0.270 0.467 0.484 0.407
+FLUX.1-dev 0.320 0.501 0.494 0.438
+Qwen-Image 0.330 0.501 0.518 0.450
+DALL-E 3 0.346 0.554 0.553 0.484
+Gemini 0.410 0.579 0.583 0.524
+Image-1 (low quality) 0.420 0.585 0.589 0.531
+Image-1 (high quality) 0.460 0.632 0.605 0.566
+Retriever = E5-V
+Original Query 0.340 0.465 0.486 0.430
+VisRet Variants:
+Stable Diffusion 3.5 0.410 0.550 0.547 0.502
+FLUX.1-dev 0.370 0.535 0.532 0.479
+Qwen-Image 0.410 0.542 0.548 0.500
+DALL-E 3 0.450 0.599 0.587 0.545
+Gemini 0.615 0.585 0.666 0.589
+Image-1 (low quality) 0.520 0.629 0.612 0.587
+Image-1 (high quality) 0.560 0.643 0.622 0.608
+Table 5: nDCG@k retrieval results on Visual-RAG-ME for different T2I generation models used in VisRet.
+We report N@1, N@10, N@30, and their average (Avg) for two retrievers (CLIP and E5-V).
+Best numbers within each retriever are boldfaced.
+Gemini stands for the Gemini-2.5-Flash-Image-Preview model.
+Knowledge Visual-RAG Visual-RAG-ME
+# images GPT-4o-mini GPT-4o GPT-4.1 # images GPT-4o-mini GPT-4o GPT-4.1
+Model Knowledge Only 0 0.385 0.485 0.492 0 0.410 0.510 0.470
+Direct T2I Retrieval 1 0.409 0.474 0.515 2 0.490 0.590 0.610
+10 0.460 0.518 0.571 10 0.480 0.630 0.650
+Visualize-then-Retrieve 1 0.418 0.492 0.572 2 0.530 0.640 0.620
+10 0.464 0.538 0.567 10 0.550 0.700 0.710
+Table 6: VQA performance comparison using different LVLMs as instruction generators for VisRet and query rephrase models.
+CLIP is used as the retriever.
+Boldfaced numbers indicate the best in each column.
+Knowledge Visual-RAG Visual-RAG-ME
+# images GPT-4o-mini GPT-4o GPT-4.1 # images GPT-4o-mini GPT-4o GPT-4.1
+Model Knowledge Only 0 0.385 0.485 0.492 0 0.410 0.510 0.470
+Generated Image (Image-1) 1 0.431 0.425 0.444 2 0.590 0.580 0.800
+Visualize-then-Retrieve 1 0.418 0.492 0.572 2 0.530 0.640 0.620
+10 0.464 0.538 0.567 10 0.550 0.700 0.710
+Table 7: VQA performance comparison using different knowledge contexts on Visual-RAG and Visual-RAG-ME. CLIP is used as the retriever. Boldfaced numbers indicate the best in each column.
+6.2 T2I Generation Model Choice
+How strong does the T2I generation model need to be for VisRet to work well?
+We compare the default T2I generation model (Image-1 with high quality setting) with several other commonly used strong T2I generation models:
+  • Stable Diffusion 3 (Esser et al., 2024).
+  • FLUX-1.dev (Black Forest Labs, 2024).
+  • Qwen-Image (Wu et al., 2025a).
+  • DALL-E 3 API (OpenAI, 2023a).
+  • Gemini-2.5-Flash-Image-Preview API (Fortin et al., 2025).
+  • Image-1 API with the low quality setting.
+Table˜5 shows the results with GPT-4o as the T2I instruction generation model and CLIP as the retriever model.
+We observe that while all the models can improve over cross-modality retrieval to some extent in the VisRet pipeline, their effectiveness increases as the image quality retrieves.
+The best performance is achieved by the latest proprietary models Image-1 and Gemini-2.5-Flash-Image-Preview.
+Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet.
+As further T2I generation methods improve, we anticipate that building more cost-efficient versions of VisRet is viable and promising.
+6.3 Downstream VQA Model Choice
+While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities?
+In Table˜6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14).
+Overall, we observe similar trends as those presented in Figure˜2.
+Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former.
+These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
+6.4 Image Queries as Knowledge
+As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet.
+A natural question is whether it is still necessary to perform retrieval instead of directly using the generated images as the knowledge?
+In Table˜7, we compare the performance of using a single image as the context with VisRet.
+Overall, we observe mixed results.
+For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern.
+For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefer the generated image over top-1 retrieval (and even top-10 retrieval).
+However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than when using the generated image.
+Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-ME and cannot be fully replaced by pure T2I generation at this stage.
+An important further direction is to combine image generation and image retrieval to improve the quality of the retrieved knowledge.
 7 Conclusion
-This work introduces VisRet, a framework that visualizes text queries to enable more accurate T2I retrieval.
-By operating entirely in the visual domain during retrieval, VisRet addresses key limitations of cross-modal embedding alignment.
-Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across three benchmarking datasets and two retrievers.
-The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multi-modal systems.
+This work introduces VisRet, a novel framework that visualizes text queries to reframe cross-modality T2I retrieval as within-modality retrieval.
+By operating entirely in the visual domain, VisRet addresses key limitations of cross-modal embedding alignment.
+Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across four benchmark datasets and two retrievers.
+The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multimodal systems.
 Ethics Statement
@@
 Although the goal of this paper is to introduce techniques to improve the text-to-image retrieval performance, the new approach could create new social risks.
-Specifically, in addition to the embedding model, our approach involves two neural networks: an LLM and a T2I generation model.
-It is possible for these large models to bring in new social bias in generating the visualized query and thus bias the retrieval results.
-For instance, when depicting certain scenes of social activity, T2I models could reinforce stereotypical social roles (Wan and Chang, 2024; Wan et al., 2024a).
+Specifically, in addition to the neural embedding model, our approach involves two neural models: an LLM and a T2I generation model.
+It is possible for these large models to bring in new social bias in generating the visualize query and thus bias the retrieval results.
+For instance, when depicting certain scenes of social activity, the models could reinforce stereotypical social roles.
 We urge practitioners to implement model debiasing and bias detection measure when deploying our proposed T2I retrieval method in real-world applications.
-Additionally, previous work has revealed problems with faithfulness and factuality in T2I models (Wan et al., 2024b; tse Huang et al., 2025; Wan and Chang, 2025).
-We also encourage future studies to explore solutions to similar trustworthy issue with multi-modal generative models.
 Artifact Release
@@
 Following Visual-RAG, we will release our Visual-RAG-ME annotations under CC BY-NC 4.0 license.
+AI Assistant Use
+AI assistants, specifically ChatGPT, are used only for revising the paper draft, fixing grammar mistakes, and improving the outlook of the figures.
 Acknowledgment
@@
 Appendix A Benchmark Data Details
-In this section, we present the details of Visual-RAG data processing and Visual-RAG-ME annotation.
-Then, we provide the details of how we processed INQUIRE-Rerenk-Hard.
-Finally, we document the dataset statistics and the baseline T2I retrieval performance of CLIP (Radford et al., 2021) and E5-V (Jiang et al., 2024) in Table 3.
-A.1 Visual-RAG and Visual-RAG-ME
-Visual-RAG (Wu et al., 2025) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^1] https://github.com/visual-rag/visual-rag.
+In this section, we present the details of data processing for Visual-RAG, INQUIRE-Rerenk-Hard, and COCO-Hard.
+A.1 Visual-RAG
+Visual-RAG (Wu et al., 2025b) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^2] https://github.com/visual-rag/visual-rag.
 To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test set.
 We were able to identify all annotated images in Visual-RAG except a single image due to a likely path error.
-We annotate Visual-RAG-ME as an extension of Visual-RAG to visual question answering in the multi-entity setting.
-Concretely, Visual-RAG-ME reuses the visual features queried in Visual-RAG and constructs questions that compares those features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels.
-The Visual-RAG-ME annotation pipeline consists of three steps: second entity identification; query composition and filtering; retrieval label annotation and balancing.
-We next describe the steps in detail.
-Second Entity Identification
-The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions could be constructed.
-For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities that have the closest full scientific names.
-Query Composition and Filtering
-In this step, we (the authors) manually traverse all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question.
-A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG.
-The question we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism feature a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.).
-We were able to construct 82 multi-entity questions after this step.
-Next, we perform a round of filtering to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions can cover overly similar topics.
-After the filtering step, we ended up with 50 high quality multi-entity queries.
-Retrieval Label Annotation and Balancing
-Finally, for each question, we collect images of the second entity from iNaturalist and annotate their retrieval label.
-A positive label is assigned only if the image clearly displays the feature required to answer the question.
-For some of the questions, we find that a large number of positive images exist in the iNaturalist database.
-We therefore implement a filtering step where at most 50 positive images are kept for each entity.
-Table 8 shows two examples with their questions and ground-truth images.
-In Table 3, we present the basic statistics of Visual-RAG-ME.
-While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
-VR VR-ME IR-Hard
-# Queries 391 50 59
-||||Query|||| (word count) 18.5 25.1 6.0
-# Images (per entity) 264 263 100
-# Positives (per entity) 14.3 20.8 12.5
-CLIP Recall@1 0.210 0.220 0.000
-E5-V Recall@1 0.240 0.340 0.000
-Table 3: Dataset statistics and baseline performance. VR denotes Visual-RAG and IR denotes INQUIRE-Rerank.
-Method R@1 R@10 R@30 N@1 N@10 N@30
-Baselines
-Original Query 0.210 0.583 0.737 0.210 0.355 0.385
-LLM Rephrase 0.238 0.586 0.737 0.238 0.360 0.395
-VisRet
-DALL-E 3 0.169 0.581 0.757 0.169 0.344 0.376
-Stable Diffusion 3 0.166 0.517 0.691 0.166 0.319 0.349
-Image-1 (low quality) 0.243 0.611 0.760 0.243 0.397 0.415
-Image-1 (high quality) 0.251 0.645 0.793 0.251 0.431 0.438
-Table 4: T2I Retrieval performance across different T2I generation models used for VisRet.
-We use GPT-4o to generate the T2I instruction and CLIP as the retriever.
-R = Recall.
-N = NDCG.
-The best results are boldfaced.
-Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
-Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
-[…]
-Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
-GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
-[…]
-Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
-GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
-Table 5: Retrieval performance on Visual-RAG of with CLIP retriever, using LLMs as T2I instruction generator for VisRet. R = Recall. N = NDCG. The best results are boldfaced.
 A.2 INQUIRE-Rerank-Hard
@@
 After manually inspecting the data, we found that for a lot of instances, the negative images are not challenging enough and it is often very straightforward to identify the ground truth images.
-To highlight the challenging questions, we therefore filtered out the questions on which either CLIP and E5-V can achieve a perfect Recall@1.
-Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organism themselves, with more challenging confounder negative images.
-Knowledge Visual-RAG Visual-RAG-ME
-# images GPT-4o-mini GPT-4o GPT-4.1 # images GPT-4o-mini GPT-4o GPT-4.1
-Model Knowledge Only 0 38.49 48.47 49.23 0 41.00 51.00 47.00
-[…]
-10 46.04 51.79 57.06 10 48.00 63.00 65.00
-[…]
-10 46.42 53.84 56.65 10 55.00 70.00 71.00
-Table 6: VQA performance comparison using different LVLMs as instruction generators for VisRet and query rephrase models.
-CLIP is used as the retriever.
-Boldfaced numbers indicate the best in each column.
-Knowledge Visual-RAG Visual-RAG-ME
-# images GPT-4o-mini GPT-4o GPT-4.1 # images GPT-4o-mini GPT-4o GPT-4.1
-Model Knowledge Only 0 38.49 48.47 49.23 0 41.00 51.00 47.00
-[…]
-[…]
-10 46.42 53.84 56.65 10 55.00 70.00 71.00
-Table 7: VQA performance comparison using different knowledge contexts on Visual-RAG and Visual-RAG-ME. CLIP is used as the retriever. Boldfaced numbers indicate the best in each column.
-Appendix B VisRet: Further Analyses
-In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I generation Model, T2I Instruction LLM, the downstream VQA LVLM reader.
-Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
-B.1 T2I Generation Model Choice
-How strong does the T2I generation model need to be for VisRet to work well?
-We compare the default T2I generation model (Image-1 with high quality setting) with three other models: DALL-E 3 (OpenAI, 2023a), Stable Diffusion 3 (Esser et al., 2024), and Image-1 with the low generation quality setting.
-Table 4 shows the results with GPT-4o as T2I instruction generation model and CLIP as the retriever model.
-Interestingly, compared to the cross-modality retrieval baselines, we find that DALL-E 3 and Stable Diffusion 3 do not provide significant performance improvements, while Image-1 low quality clearly and consistently improve the performance.
-The best performance is achieved by the newest and the most expensive Image-1 high quality setting.
-Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet.
-As further T2I generation methods improve, we anticipate that building more cost-efficient version of VisRet is a viable and promising further direction.
-B.2 T2I Instruction LLM Choice
-Does VisRet work well with other LLMs as the T2I instruction generator?
-In Table 5, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024).
-Overall, we observe promising results.
-For all the three LLMs, using them to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query.
-While more expensive LLMs achieve a high performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
-B.3 Downstream VQA LVLM Choice
-While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities?
-In Table 6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14).
-Overall, we observe similar trends as those presented in Figure 2.
-Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former.
-These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
-B.4 Image Queries as Knowledge
-As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet.
-It is a natural question then, that it is still necessary to perform retrieval instead of directly using the generated images as the knowledge?
-In Table 7, we compare the performance of using a single image as the context with VisRet.
-Overall, we observe a mixed result.
-For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern.
-For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefers the generated image over top-1 retrieval (and even top-10 retrieval).
-However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than using the generated image.
-Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-MR and cannot be fully replaced by pure T2I generation at this stage.
-It is an important future work to further combine image generation and image retrieval to improve the quality of the retrieved knowledge.
-B.5 Further Qualitative Studies
-In Table 8, we additionally provide some qualitative examples of how VisRet successfully improves T2I retrieval performance.
-As can be observed in top-ranked examples in text query-only retrieval, these retrieved images often fall short in correctly conveying visual semantics information such as angle, body part depicted, form of the subject depicted, visual distance of the subject, and so on.
-In contrast, VisRet is able to represent the inferred nuanced visual semantics in the image generation step first, then utilize this within-modality semantic-rich image query to obtain precise visual knowledge related to the VQA task.
-Appendix C Implementation Details
+To highlight the challenging questions, we therefore filtered out the questions on which either CLIP or E5-V can achieve a perfect Recall@1.
+Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organisms themselves, with more challenging confounder negative images.
+A.3 COCO-Hard
+We derive COCO-Hard from the publicly released validation set of Microsoft COCO 2017 (Chen et al., 2015).
+The original validation set contained 5000 queries, and we directly use the ground truth images themselves to construct the image corpus.
+In a pilot study, we found that CLIP is able to achieve 0.424 Recall@1 and 0.756 Recall@10, while E5-V has 0.594 Recall@1 and 0.892 Recall@10.
+A manual analysis shows that the queries of most of the examples in the original COCO dataset are short phrases centered on a single entity.
+To increase the task’s difficulty, we rank the examples in reverse order by nDCG@30 and retained the hardest 500 instances.
+We find these hard queries often describe complex spatial relationships among multiple entities.
+Appendix B Implementation Details
 In this section, we present the implementation details of VisRet and baselines.
@@
 To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in it.
-The prompts for three benchmarking datasets are shown in Figure 3, Figure 4, and Figure 5, respectively.
-Then, we wrap the rephrased query with a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation.
+The prompts for three benchmarking datasets are shown in Figure˜3, Figure˜4, and Figure˜5, respectively.
+Then, we wrap the rephrased query in a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation.
 We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high".
-For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a signle instruction already results in images generated with a high level of diversity.
+For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a single instruction already results in images generated with a high level of diversity.
 Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
@@
 Concretely, the model is asked to always extract visual information, perform reasoning, and conclude its reasoning with self-verification.
-We show the detailed prompt in Figure 6 and Figure 7.
+We show the detailed prompt in Figure˜6 and Figure˜7.
+VQA Evaluation
+For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025b), as shown in Figure˜8.
+Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering.
+We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
 Baselines
 For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important feature that the query is seeking.
-At the early stage of the prompt, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction.
+At the early stage of the project, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction.
 Therefore, we report the results with the same prompt for the final version for the paper.
-VQA Evaluation
-For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025), as shown in Figure 8.
-Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering.
-We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
-Dataset Question Ground Truth Baseline Generated Image VisRet
+For the Corpus Captioning baseline, we use the Salesforce/blip-image-captioning-large model accessed at huggingface.
+We follow the recommended parameters to run inference on local GPUs, except setting the minimum output length to 10 tokens to encourage a more detailed caption.
+For the VISA Reranking baseline, we followed the original paper’s public implementation [^6] https://github.com/XLearning-SCU/2025-ICML-VISA, using the sample prompt except changing the model to Qwen2.5-VL-72B-Instruct which we empirically find better performance.
+For the Captioning Reranking baseline, we simplify the VISA pipeline by directly prompting for a caption for each of the top retrieved images using the following prompt:
+"Write a caption that describes the appearance and posture of the main subject.
+The caption should be concise but informative so that it can be used to retrieve the images from images of the same subject in different postures, states, or environments."
+Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
 Visual-RAG What is the color of the head of larva Urania Swallowtail Moth (scientific name: Urania fulgens)? [Uncaptioned image] Rank:96, nDCG@10: 0.00 [Uncaptioned image] Rank:1, nDCG@10: 0.66
@@
 Are there visually distinctive scales on feet of Azure Tit (scientific name: Cyanistes cyanus)? [Uncaptioned image] Rank:26, nDCG@10: 0.40 [Uncaptioned image] Rank:2, nDCG@10: 0.76
-INQUIRE A male and female cardinal sharing food [Uncaptioned image] Rank:12, NDCG@10: 0.00 [Uncaptioned image] Rank:1, NDCG@10: 1.00
-Mexican grass-carrying wasp visiting a purple flower [Uncaptioned image] Rank:45, NDCG@10: 0.00 [Uncaptioned image] Rank:4, NDCG@10: 0.39
+INQUIRE Mexican grass-carrying wasp visiting a purple flower [Uncaptioned image] Rank:45, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.39
 great golden digger wasp carrying an orthopteron [Uncaptioned image] Rank:12, nDCG@10: 0.45 [Uncaptioned image] Rank:2, nDCG@10: 0.83
@@
 [Uncaptioned image] Entity: theona checkerspot Rank:120, nDCG@10: 0.52 [Uncaptioned image] Entity: theona checkerspot Rank:3, nDCG@10: 0.85
+COCO A den with a large coffee table, couch and a television. [Uncaptioned image] Rank:61, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.3937
+A computer workstation with a laptop and a desktop computer. [Uncaptioned image] Rank:37, nDCG@10: 0.00 [Uncaptioned image] Rank:1, nDCG@10: 1.0
+A residential bathroom with sink, tub, and toilet setting in it. [Uncaptioned image] Rank:54, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.39
 Table 8: Additional qualitative results on the three benchmarking datasets.
@@
 Original query: {question} Rephrased query:
-Figure 5: Prompt for instructing an LLM to generate the T2I generation instruction for INQUIRE-Rerank questions.
+Figure 5: Prompt for instructing an LLM to generate the T2I generation instruction for INQUIRE-Rerank-Hard and COCO-Hard questions.
 […]
@@
 Figure 8: Prompt for the LLM VQA judge used for Visual-RAG and Visual-RAG-ME.
+Appendix C Further Qualitative Examples
+In Table˜8, we present additional qualitative examples on the four evaluation datasets.
+Consistent with quantitative findings, VisRet reduces the embedding’s workload by expressing the important details in a visualization close to the retrieval target, resulting in much better performance in terms of ground truth rank and nDCG@10.

### v3

@@
 Abstract
-Text-to-image retrieval (T2I retrieval) remains challenging because cross-modal embeddings often behave as bags of concepts and underrepresent structured visual relationships such as pose and viewpoint.
-We propose Visualize-then-Retrieve (VisRet), a new paradigm for T2I retrieval that mitigates this limitation of cross-modal similarity alignment.
-VisRet first projects textual queries into the image modality via T2I generation.
-Then, it performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features.
+Text-to-image retrieval (T2I retrieval) remains challenging because cross-modal embeddings often behave as bags of concepts, underrepresenting structured visual relationships such as pose and viewpoint.
+We propose Visualize-then-Retrieve (VisRet), a retrieval paradigm that mitigates this limitation of cross-modal similarity alignment.
+VisRet first projects textual queries into the image modality via T2I generation, then performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features.
 Across four benchmarks (Visual-RAG, INQUIRE-Rerank, Microsoft COCO, and our new Visual-RAG-ME featuring multi-entity comparisons), VisRet substantially outperforms cross-modal similarity matching and baselines that recast T2I retrieval as text-to-text similarity matching, improving nDCG@30 by 0.125 on average with CLIP as the retriever and by 0.121 with E5-V.
@@
 Ablation studies show compatibility with different T2I instruction LLMs, T2I generation models, and downstream LLMs.
-VisRet provides a practical and principled path that energizes further advances in vision-language retrieval.
+VisRet provides a simple yet effective perspective for advancing in text-image retrieval.
 Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
@@
 1 Introduction
-Text-to-Image (T2I) retrieval is the task of selecting the most relevant images from a visual corpus based on a textual query.
-It plays a crucial role in enabling knowledge-intensive applications that require supporting textual inputs with rich visual content (Chen et al., 2022; Wang et al., 2023; Sheynin et al., 2023; Braun et al., 2024).
-A common approach to T2I retrieval is to embed both the text query and image candidates into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014).
-However, obtaining accurate similarity rankings that capture fine-grained semantic alignments between text and image remains a long-standing challenge.
-Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts” for similarity matching, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023).
-For instance, Figure˜1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer.
-While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (upward-facing shot).
-To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025).
-Nevertheless, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
+Text-to-Image (T2I) retrieval selects the most relevant images from a visual corpus given a textual query.
+It plays a crucial role in knowledge-intensive applications that augment textual inputs with rich visual evidence (DBLP:conf/emnlp/ChenHCVC22; DBLP:conf/acl/Wang0LLCJHXG23; DBLP:conf/iclr/SheyninAPSGNT23; DBLP:journals/corr/abs-2412-10510).
+A common approach embeds both the text query and image candidates into a shared representation space and ranks candidates by similarity (frome2013devise; DBLP:journals/corr/KirosSZ14).
+However, producing rankings that reflect fine-grained alignments between text and image remains a long-standing challenge.
+Prior studies show that cross-modal embeddings often behave like “bags of concepts” for similarity matching and fail to capture structured relationships among visual elements (DBLP:conf/iclr/Yuksekgonul0KJ023; DBLP:conf/emnlp/KamathHC23).
+For example, Figure˜1 presents a query that requires images of an entity (a Barnacle Goose) in a specific posture (wings unfolded).
+While the embedding model matches the entity type, it struggles to recognize subtler visual-spatial features such as the wing pose and the camera perspective (an upward-facing shot).
+To address these limitations, existing work improves embedding quality (DBLP:conf/icml/RadfordKHRGASAM21; DBLP:journals/tmlr/YuWVYSW22), performs query reformulation (levy2023chatting), or applies multi-stage reranking pipelines (DBLP:journals/tmlr/0002STG24; DBLP:conf/aaai/0001B00KZGL25).
+Nevertheless, these strategies remain constrained by the intrinsic difficulty of cross-modal similarity alignment because they cannot bypass the text-to-image similarity search stage.
 Figure 1: An overview of VisRet.
 Compared to the traditional T2I retrieval pipeline, VisRet first projects the text query into the image modality via T2I generation and then performs within-modality retrieval.
-We propose Visualize-then-Retrieve (VisRet), a novel retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval.
-Concretely, VisRet first utilizes a text-to-image generation (T2I generation) model to visualize the textual query, highlighting the key desired visual-spatial features.
-Then, the visualized query is used to perform direct image-to-image retrieval.
-Compared to prior methods, VisRet offers two key advantages.
-First, visualizations provide a more expressive and intuitive medium for representing multiple compositional concepts, including entities, poses, and spatial relations.
-These concepts are often difficult to convey through text alone, and attempting to encode them exhaustively in a text query can impair cross-modal matching performance due to limitations in embedding quality.
-In contrast, as shown in Figure˜1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time.
-Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in unimodal retrieval (Koishigarina et al., 2025).
-To investigate the effectiveness of VisRet, we conduct exhaustive experiments on two commonly used cross-modal embedding models using four benchmarks: Visual-RAG (Wu et al., 2025b), INQUIRE-Rerank (Vendrow et al., 2024), Microsoft COCO (Chen et al., 2015), and Visual-RAG-ME, a new benchmark we introduce that features challenging visual feature comparison questions across multiple entities.
-Results show that VisRet substantially outperforms baseline cross-modal similarity matching as well as three strong baseline methods that casts T2I retrieval to text-to-text similarity matching (section˜5.2).
-When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the baselines by 0.125 higher nDCG@30, averaged over four benchmarks.
-With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 0.121.
-Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section˜5.3).
-It improves T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval setting and 3.9% and 11.1% in top-10 retrieval setting.
-Further analyses show that VisRet is compatible with different T2I instruction generation and downstream question answering LLMs.
-In addition, the retrieval performance consistently improves with stronger T2I models.
-Finally, although they are effective in retrieving visual knowledge, the generated image queries themselves are generally insufficient as standalone knowledge to answer downstream questions.
+We propose Visualize-then-Retrieve (VisRet), a retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval.
+VisRet first uses a text-to-image generation (T2I generation) model to visualize the textual query, with the goal of making key visual-spatial requirements explicit.
+The resulting visualization is then used as a query for direct image-to-image retrieval.
+Compared to prior methods, VisRet offers two advantages.
+First, visualizations provide a more expressive and intuitive medium for representing compositional concepts such as entities, poses, and spatial relations.
+These requirements can be difficult to communicate through text alone, and encoding them exhaustively in a text query can hurt cross-modal matching due to limitations in embedding quality.
+In contrast, as shown in Figure˜1, the visualized query can depict the desired entity, posture, and camera angle at the same time.
+Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and leverages the stronger capacity of these retrievers in unimodal retrieval (DBLP:journals/corr/abs-2502-03566).
+To evaluate VisRet, we conduct experiments with two widely used cross-modal embedding models across four benchmarks: Visual-RAG (visual-rag), INQUIRE-Rerank (DBLP:conf/nips/VendrowPSBJABH24), Microsoft COCO (mscoco), and Visual-RAG-ME, a new benchmark we introduce that features challenging visual feature comparison questions across multiple entities.
+VisRet substantially outperforms baseline cross-modal similarity matching and three strong baseline methods that cast T2I retrieval as text-to-text similarity matching (section˜5.2).
+When CLIP (DBLP:conf/icml/RadfordKHRGASAM21) is used as the retriever, VisRet improves nDCG@30 by 0.125 on average across the four benchmarks.
+With E5-V (DBLP:journals/corr/abs-2407-12580) as the retriever, the improvement is 0.121.
+VisRet also improves downstream performance in retrieval-augmented generation (RAG) settings (section˜5.3), increasing T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval and by 3.9% and 11.1% in top-10 retrieval.
+Ablation studies and analyses show that VisRet is robust to the choice of instruction-generation LLM, retriever, and downstream reader LVLM, while performance depends more strongly on the T2I generation model, with stronger generators yielding larger gains.
+We further analyze repeated runs and find that aggregating multiple visualizations is beneficial, while using a single visualization only slightly reduces overall performance.
+Finally, although visualized queries are effective for retrieval, they are generally insufficient as standalone knowledge to answer downstream questions.
+Our code and the new Visual-RAG-ME benchmark will be released publicly to facilitate future research.
 2 Related Work
 T2I Retrieval Benchmarks
-Early T2I retrieval benchmarks evaluate the ability to identify images based on their paired human-written captions.
-These datasets span multiple domains and include widely used benchmarks such as COCO (Chen et al., 2015), Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017).
-As multimodal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings.
-These newer datasets, such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025b), and MRAG-Bench (Hu et al., 2024), shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions.
-These tasks challenge T2I retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines.
-Further extending this stream of work, our Visual-RAG-ME benchmarks T2I retrieval for reasoning over the visual features across multiple entities.
+Early T2I retrieval benchmarks evaluate whether a system can retrieve an image given its paired human-written caption.
+These datasets span multiple domains and include widely used benchmarks such as COCO (mscoco), Flickr8K (DBLP:journals/jair/HodoshYH13), Flickr30K (DBLP:journals/tacl/YoungLHH14), and Fashion200K (han2017automatic).
+As multimodal embedding models have matured, newer benchmarks have been introduced to assess retrieval in knowledge-intensive settings.
+Datasets such as WebQA (DBLP:conf/cvpr/ChangCNGSB22), INQUIRE (DBLP:conf/nips/VendrowPSBJABH24), Visual-RAG (visual-rag), and MRAG-Bench (DBLP:journals/corr/abs-2410-08182) shift the focus from caption matching to retrieving images that provide the information needed to answer knowledge-intensive questions.
+These tasks evaluate T2I retrieval systems as components of retrieval-augmented generation (RAG) pipelines and emphasize support for downstream question answering.
+Extending this line of work, Visual-RAG-ME benchmarks T2I retrieval for reasoning about visual features shared across multiple entities.
 T2I Retrieval Methods
-Extensive research has focused on improving text-to-image retrieval from different perspectives.
-First, a number of works aim to train better multimodal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022).
-Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025; Ding et al., 2025).
-Finally, a recent line of work introduces generative image retrieval (Li et al., 2024; Qu et al., 2025), which trains a generative model to directly memorize an index of the image corpus.
-Different from these existing approaches, VisRet expands the query semantics by directly visualizing it in the image space, thereby alleviating the workload of cross-modal retrieval.
-In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever and pre-computed image embedding indices.
+A large body of work improves T2I retrieval from several angles.
+First, many approaches aim to train stronger multimodal embeddings through improved objectives and data mixtures (DBLP:conf/bmvc/FaghriFKF18; DBLP:conf/icml/RadfordKHRGASAM21; DBLP:journals/tmlr/YuWVYSW22; DBLP:conf/icml/0001LXH22).
+Other studies target specific stages of the pipeline, including textual query expansion (levy2023chatting; DBLP:conf/acl/LeeYPYY24) and reranking (DBLP:journals/tmlr/0002STG24; DBLP:conf/aaai/0001B00KZGL25; ding2025visual).
+Finally, a recent line of work explores generative image retrieval (DBLP:conf/acl/LiWQNLC24; qu2025tigerunifyingtexttoimagegeneration), which trains a generative model to directly memorize an index of the image corpus.
+In contrast, VisRet expands query semantics by visualizing the query in the image space, which reduces the burden on cross-modal retrieval.
+It is also training-free and plug-and-play, accommodating off-the-shelf retrievers or pre-computed image embedding indices.
 3 Approach
 3.1 Problem Formulation
-Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics of $q$.
-Our paper further considers the task of Visual Question Answering (VQA), where the query is a knowledge-seeking question, with an expected answer $a$.
-In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multimodal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$.
-Then, a large vision-language model (LVLM) $\mathcal{M}$ directly generates the answer based on the question and the retrieval results $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
+Given a textual query $q$ and an image corpus $\mathcal{I}$, Text-to-Image Retrieval aims to retrieve $n\geq 1$ images $y_{1},\ldots,y_{n}\in\mathcal{I}$ that best match the semantics of $q$.
+We further consider Visual Question Answering (VQA), where the query is a knowledge-seeking question with an expected answer $a$.
+We consider a basic retrieval-augmented generation (RAG) pipeline for VQA.
+A multimodal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},\ldots,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$.
+A large vision-language model (LVLM) $\mathcal{M}$ then generates an answer conditioned on the question and retrieval results, i.e., $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
 3.2 Visualize-then-Retrieve
-We introduce Visualize-then-Retrieve (VisRet), a two-stage T2I retrieval pipeline that bridges the modality gap through modality projection.
-Figure˜1 illustrates the pipeline with an intuitive example.
+We introduce Visualize-then-Retrieve (VisRet), a two-stage T2I retrieval pipeline that bridges the modality gap through modality projection. Figure˜1 illustrates the pipeline with an example.
 Modality Projection
-The first stage of VisRet leverages a T2I generation system $\mathcal{T}$ to directly generate $m$ visualizations $\{v_{1},...,v_{m}\}\equiv\mathcal{T}(q)$.
-Concretely, based on the original query, an LLM first operates in the text space to draft T2I instruction $q^{\prime}$, describing the images that likely satisfy the feature requirements implied in $q$.
-Next, we use an existing T2I generation method, such as Stable Diffusion (Esser et al., 2024), to project $q^{\prime}$ into a list of image queries $\{v_{1},...,v_{m}\},v_{i}\in\mathcal{I}$.
-To generate diverse images, randomness can be injected either into $q^{\prime}$ or into the T2I generation.
-In this work, we sample $m$ times with the same $q^{\prime}$.
+In the first stage, VisRet uses a T2I generation system $\mathcal{T}$ to synthesize $m$ visualizations $\{v_{1},\ldots,v_{m}\}\equiv\mathcal{T}(q)$.
+Concretely, given the original query, an LLM first operates in the text space to draft a T2I instruction $q^{\prime}$ that describes images likely to satisfy the feature requirements implied by $q$.
+We then use an existing T2I generation method, such as Stable Diffusion (DBLP:conf/icml/EsserKBEMSLLSBP24), to project $q^{\prime}$ into a set of image queries $\{v_{1},\ldots,v_{m}\}$, where each $v_{i}$ is a synthesized image.
+To encourage diversity, randomness can be introduced either in the instruction drafting step or within the T2I generation process.
+In this work, we sample $m$ times using the same $q^{\prime}$.
 Within-Modality Retrieval
-In the second stage, VisRet performs retrieval exclusively within the image modality.
-First, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
-$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},...,r_{k}^{(i)}],$
-Then, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) to aggregate the $m$ retrieval results. RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
+In the second stage, VisRet performs retrieval entirely within the image modality.
+Each synthesized image $v_{i}\in\{v_{1},\ldots,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
+$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},\ldots,r_{k}^{(i)}].$
+We then apply Reciprocal Rank Fusion (RRF) (DBLP:conf/sigir/CormackCB09) to aggregate the $m$ ranked lists. RRF assigns each candidate image $r$ a fusion score based on its rank across the $m$ lists:
 $\text{score}_{\text{RRF}}(r)=\sum_{i=1}^{m}\frac{1}{\lambda+\text{rank}_{i}(r)},$
-where $\text{rank}_{i}(r)$ is the rank position of image $r$ in list $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items.
-The final top-$k$ retrieval result is obtained by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
+where $\text{rank}_{i}(r)$ is the rank position of image $r$ in $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items.
+The final top-$k$ results are obtained by selecting images with the highest $\text{score}_{\text{RRF}}(r)$.
 4 Visual-RAG-ME Benchmark
-To facilitate the design of intelligent T2I retrieval systems, we introduce a multi-entity question answering and T2I retrieval dataset Visual-RAG-ME as an extension of Visual-RAG.
-Concretely, Visual-RAG-ME constructs questions that compare the visual features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels.
-The Visual-RAG-ME annotation pipeline consists of three steps as described below.
+To support the development of stronger T2I retrieval systems, we introduce Visual-RAG-ME, a multi-entity question answering and T2I retrieval dataset that extends Visual-RAG.
+Visual-RAG-ME constructs questions that compare visual features between the organism in a Visual-RAG query and another similar entity, and we manually annotate a new set of retrieval labels for the second entity.
+The Visual-RAG-ME annotation pipeline consists of the following three steps.
 Second Entity Identification
-The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions can be constructed.
-For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities with the most similar scientific names.
+This step identifies potential entities that are biologically close to the original Visual-RAG entities, enabling plausible and challenging comparison questions.
+For each entity in Visual-RAG, we use BM25 (10.1007/978-1-4471-2099-5_24) to retrieve up to ten entities with the most similar scientific names.
 Query Composition and Filtering
-In this step, we (the authors) manually review all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question.
-A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG.
-The questions we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism features a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.).
-We were able to construct 82 multi-entity questions after this step.
-Then, a round of filtering is performed to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions that cover overly similar topics.
+We (the authors) manually review all 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question.
+We construct a question only when we can identify images for the second entity that clearly depict the same feature as the positive images for the original entity in Visual-RAG.
+The questions we compose typically use a comparison format that asks whether the two organisms share a feature or which organism exhibits a more extreme stylistic feature (e.g., lighter coloration, smoother surface).
+This process yields 82 multi-entity questions.
+We then filter out (1) questions answerable by both GPT-4o and GPT-4o-mini without image information and (2) excessive questions covering the same topic.
 The final Visual-RAG-ME contains 50 high-quality multi-entity queries.
 Retrieval Label Annotation and Balancing
-For each question, we collect images of the second entity from the iNaturalist 2021 dataset (Horn et al., 2021) and annotate their retrieval labels.
-A positive label is assigned only if the image clearly displays the feature required to answer the question.
-For some of the questions, we find that a large number of positive images exist in the iNaturalist database.
-We therefore implement a filtering step where at most 50 positive images are kept for each entity.
-Table˜3 and Appendix Table˜8 show examples with their questions and ground-truth images.
-In Table˜1, we present the basic statistics of Visual-RAG-ME.
-While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
+For each question, we collect images of the second entity from the iNaturalist 2021 dataset (DBLP:conf/cvpr/HornCBWBA21) and annotate retrieval labels.
+We assign a positive label only when the image clearly displays the feature required to answer the question.
+For some questions, many positive images exist in iNaturalist, so we apply a filtering step that keeps at most 50 positive images for each entity.
+Table˜5 and Appendix Table˜10 show example questions and their ground-truth images.
 VR VR-ME IR-Hard $C$-Hard
@@
 Table 1: Dataset statistics and baseline performance. VR = Visual-RAG, IR = INQUIRE-Rerank, and $C$ = Microsoft COCO 2017.
+Table˜1 reports the basic statistics of Visual-RAG-ME.
+While it contains slightly more positives per entity than Visual-RAG, it remains challenging.
+Both CLIP and E5-V achieve low Recall@1 due to the lengthy and knowledge-intensive queries.
+5 Main Results
 Retrieval Method Visual-RAG Visual-RAG-ME INQUIRE-Rerank-Hard COCO-Hard
@@
 VisRet 0.251 0.431 0.438 0.460 0.632 0.605 0.170 0.452 0.455 0.034 0.072 0.108
-VisRet - multi-image 0.246 0.414 0.421 0.480 0.629 0.605 0.237 0.428 0.469 0.038 0.069 0.106
 Retriever = E5-V
@@
 VisRet 0.299 0.452 0.461 0.560 0.643 0.622 0.220 0.377 0.425 0.042 0.173 0.205
-VisRet - multi-image 0.307 0.442 0.446 0.520 0.640 0.617 0.203 0.384 0.445 0.048 0.168 0.206
 Table 2: nDCG (N@k) across four T2I retrieval benchmarks using different retrieval strategies and retrievers.
@@
 VisRet achieves the state-of-the-art across most evaluation settings, especially excelling in handling knowledge-intensive queries from Visual-RAG and Visual-RAG-ME.
-Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
-Visual-RAG Does the Mountain Tree Frog (scientific name: Hyla eximia) have any distinctive pattern on the underside of its body?
-[Uncaptioned image] Rank:49, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.39
-How many petals are on each of the Tower Mustard (scientific name: Turritis glabra)’s flowers? [Uncaptioned image] Rank:143, nDCG@10: 0.00 [Uncaptioned image] Rank:2, nDCG@10: 0.76
-Visual-RAG-ME Which one has striped primary flight feathers, Willet (scientific name: Tringa semipalmata) or Grey-tailed tattler (scientific name: Tringa brevipes)?
-[Uncaptioned image] Entity: Willet Rank:104, nDCG@10: 0.00 [Uncaptioned image] Entity: Willet Rank:1, nDCG@10: 0.72
-[Uncaptioned image] Entity: Grey-tailed tattler Rank:74, nDCG@10: 0.00 [Uncaptioned image] Entity: Grey-tailed tattler Rank:1, nDCG@10: 1.00
-INQUIRE A male and female cardinal sharing food [Uncaptioned image] Rank:12, nDCG@10: 0.00 [Uncaptioned image] Rank:1, nDCG@10: 1.00
-COCO Plate with pizza, knife and fork laying on edge of plate with two glasses next to them. [Uncaptioned image] Rank:39, nDCG@10: 0.00 [Uncaptioned image] Rank:3, nDCG@10: 0.3937
-Table 3: Examples: VisRet improves retrieval by highlighting visual features implied by the textual query.
-"Baseline" refers to cross-modal retrieval with the original query.
-CLIP embedding scores are used for this visualization.
-5 Main Results
 5.1 Experimental Setup
 Benchmarks
-We evaluate on four challenging benchmarks: (1) Visual-RAG (Wu et al., 2025b), a T2I retrieval and VQA benchmark featuring visual knowledgeintensive questions on features of natural species that are not commonly documented in text corpora.
-(2) Visual-RAG-ME, a new benchmark we introduce featuring queries that compare the same visual feature across multiple entities.
-(3) INQUIRE-Rerank (Vendrow et al., 2024), a T2I retrieval benchmark requiring accurate knowledge of species appearance and behavior.
-(4) Microsoft COCO 2017 (Chen et al., 2015), a widely used image captioning dataset, which we repurpose for caption-to-image retrieval evaluation.
-For (3) and (4), we perform additional filtering to retain the most difficult queries and call the resulting datasets INQUIRE-Rerank-Hard and COCO-Hard.
-Table˜1 shows the basic statistics of each dataset.
-We present additional details in Appendix Appendix˜A.
+We evaluate on four benchmarks: (1) Visual-RAG (visual-rag), a T2I retrieval and VQA benchmark featuring visual knowledge-intensive questions about natural species, where the relevant evidence is often not readily available in text corpora.
+(2) Visual-RAG-ME, our new benchmark that compares the same visual feature across multiple entities.
+(3) INQUIRE-Rerank (DBLP:conf/nips/VendrowPSBJABH24), which requires accurate knowledge of species appearance and behavior for retrieval.
+(4) Microsoft COCO 2017 (mscoco), a widely used image captioning dataset that we repurpose for caption-to-image retrieval evaluation.
+For (3) and (4), we further filter queries to retain the most difficult cases, yielding INQUIRE-Rerank-Hard and COCO-Hard.
+Table˜1 reports dataset statistics, and we provide additional details in Appendix Appendix˜A.
 Metrics
-For all the three benchmarks, we evaluate T2I retrieval with Recall@$k$ and nDCG@$k$ with $k$ = 1, 10, 30111As nDCG@1 is undefined, we use Recall@1 as its value..
-For Visual-RAG and Visual-RAG-ME, we additionally use an LLM judge to evaluate the end-to-end VQA accuracy, using a prompt derived from Visual-RAG (Wu et al., 2025b).
+For all four retrieval benchmarks, we report Recall@$k$ and nDCG@$k$ with $k\in\{1,10,30\}$111nDCG@1 is undefined and we use Recall@1 as its value..
+For Visual-RAG and Visual-RAG-ME, we also evaluate end-to-end VQA accuracy using an LLM judge with a prompt derived from Visual-RAG (visual-rag).
 Baselines
-Using the benchmarks, we evaluate three families of T2I retrieval methods:
+We compare three families of T2I retrieval methods:
   1. 1.
-    Text-to-Image Matching: Retrieving with the original textual query against the image index using off-the-shelf retrievers (CLIP, E5-V).
-We also compare with a baseline of applying query rewriting via an LLM to highlight the desired features and then performing standard T2I retrieval over the same image embedding index.
+    Text-to-Image Matching: Retrieve using the original textual query against the image index with off-the-shelf retrievers (CLIP, E5-V).
+We also include a query rewriting baseline, where an LLM rewrites the query to highlight desired features before performing standard T2I retrieval over the same embedding index.
   2. 2.
-    Text-to-Text Matching: Converting the images into textual surrogates for text-to-text retrieval.
+    Text-to-Text Matching: Convert images into textual surrogates and perform text-to-text retrieval avoiding cross-modal similarity matching.
 We compare with (1) captioning the entire corpus with BLIP, (2) top-k reranking with VISA, and (3) an ablated version of VISA, where we rerank top-k images using LLM-generated captions.
-We use Qwen2.5-70B-VL (Bai et al., 2025) as the model for (2) and (3) and rerank the top-30 retrieval.
   3. 3.
-    Image-to-Image Matching: We evaluate VisRet and an image ablation of VisRet that only uses a single image instead of multiple images (-multi-image).
+    Image-to-Image Matching: At the time of writing (December 2025), VisRet is the only method in this category.
 Implementation Details
-For all the retrieval experiments, we use CLIP and E5-V as the retrievers, GPT-4o (OpenAI, 2023b) as T2I instruction generation LLM, and gpt-image-1 (OpenAI, 2025) as the T2I Model to generate $m=3$ images.
-For the downstream question answering (Visual-RAG and Visual-RAG-ME), we use GPT-4o as the reader LLM.
-We present the full implementation details of VisRet and the baselines in Appendix˜B.
+For all retrieval experiments, we use CLIP and E5-V as retrievers, GPT-4o (OpenAI2023GPT4TR) as the T2I instruction generation LLM, and gpt-image-1 (openai_gpt_image_1) as the T2I model to generate $m=3$ images.
+For downstream question answering on Visual-RAG and Visual-RAG-ME, we use GPT-4o as the reader LLM.
+Full implementation details for VisRet and baselines are provided in Appendix Appendix˜B.
 5.2 VisRet Improves T2I Retrieval
 […]
-Compared with VisRet, the text-to-text retrieval baselines exhibit inconsistent performance across the benchmarks.
-Captioning the entire corpus with BLIP only works on COCO, which has the closest relationship to its training data, while it does not improve performance over cross-modal retrieval on the other benchmarks with more knowledge intensive queries.
-We hypothesize that this is caused by the high information density in images, which incurs an inevitable loss of potentially useful information during captioning.
-On the other hand, with access to cross-modal retrieval results and by directly operating on them, Captioning Reranking and VISA outperform query-agnostic corpus captioning, yet cannot outperform direct cross-modal retrieval by a large margin.
-Furthermore, these methods can only process a limited number of top-k retrieved items as their cost grows linearly with k.
-As a result, their performance is limited by the first round of retrieval, leading to near-zero performance on INQUIRE-Rerank-Hard and COCO-Hard.
-Finally, comparing VisRet with its single-image variant, we find that using only one generated image as the query slightly harms the performance but still outperforms most of the other baselines.
-The finding holds for both the lightweight CLIP embedding or the stronger E5-V embeddings, indicating the flexibility of VisRet.
-Table˜3 presents several examples to demonstrate how the visualization step successfully captures subtle visual features implied by the original text query.
-As shown in the examples, for a number of ground truth images, the main entity is either in a particular posture that reveals the required information (Visual-RAG and Visual-RAG-ME) or in a complex relationships between other entities (INQUIRE and COCO), posing a challenge to cross-modal similarity matching to recognize and satisfy these constraints.
-On the other hand, VisRet infuses these constraints naturally into the image query and significantly reduces the workload of the downstream embedding model.
+Text-to-text baselines behave less consistently across benchmarks.
+Corpus captioning with BLIP helps on COCO, which is closest to its training distribution, but does not improve over cross-modal retrieval on benchmarks with more knowledge-intensive queries.
+We hypothesize that this reflects the high information density in images and the corresponding loss of potentially useful details during captioning.
+In contrast, reranking approaches that operate on cross-modal retrieval results, including Captioning Reranking and VISA, outperform query-agnostic corpus captioning.
+However, they still do not outperform direct cross-modal retrieval by a large margin.
+These approaches are also constrained by cost, since they process only top-$k$ candidates and their cost grows linearly with $k$.
+As a result, their performance is limited by the initial retrieval stage, leading to near-zero performance on INQUIRE-Rerank-Hard and COCO-Hard.
 Figure 2: Downstream RAG-based VQA accuracy on Visual-RAG and Visual-RAG-ME with CLIP as the retriever and GPT-4o as the reader LVLM.
 5.3 VisRet Improves Downstream QA
-Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
-Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
-LLM Rewriting Llama 3.1 8B Instruct 0.238 0.563 0.737 0.238 0.365 0.385
-Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
-GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
-Visualize-then-Retrieve Llama 3.1 8B Instruct 0.243 0.606 0.780 0.243 0.405 0.428
-Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
-GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
-Table 4: Retrieval performance on Visual-RAG of with CLIP retriever, using different LLMs as T2I instruction generator for VisRet. R = Recall. N = nDCG. The best results are boldfaced.
-To assess the utility of VisRet in real-world applications, we evaluate its downstream VQA accuracy in a RAG pipeline.
-We compare three settings: (1) using only the model’s internal knowledge, (2) RAG with original text query-based retrieval and (3) RAG with VisRet.
-Figure˜2 shows the QA accuracy on Visual-RAG and Visual-RAG-ME using GPT-4o as the LVLM reader and CLIP as the retriever.
-The original query results in low-quality retrieval augmentation, even slightly harming performance on Visual-RAG in top-1 retrieval setting compared to no retrieval.
-By contrast, VisRet significantly improves QA accuracy in both top-1 and top-10 settings on both benchmarks, boosting accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME.
-Remarkably, on Visual-RAG-ME, VisRet outperforms top-10 retrieval in the original query setting with only top-1 retrieval, highlighting its high accuracy in retrieving the images that contain the required features.
-Overall, the results confirm that VisRet not only improves retrieval accuracy but also leads to tangible gains in downstream VQA performance.
-In Section˜6.3, we further demonstrate that VisRet can bring similar performance gains to other models as the VQA reader.
-In Section˜6.4, we analyze the performance of directly using the generated images as the context and find that while T2I generation improves retrieval, it still cannot replace the retrieved natural images in most cases.
+To assess VisRet in a RAG pipeline, we evaluate downstream VQA accuracy under three settings: (1) answering with only the model’s internal knowledge, (2) RAG using retrieval from the original text query, and (3) RAG using VisRet.
+Figure˜2 reports QA accuracy on Visual-RAG and Visual-RAG-ME with GPT-4o as the LVLM reader and CLIP as the retriever.
+The original query often provides low-quality retrieval augmentation, and it slightly reduces performance on Visual-RAG in the top-1 retrieval setting relative to no retrieval, suggesting that inaccurate evidence can be more harmful than no evidence.
+In contrast, VisRet improves QA accuracy in both top-1 and top-10 settings on both benchmarks, raising accuracy to 0.538 on Visual-RAG and 0.700 on Visual-RAG-ME.
+Notably, on Visual-RAG-ME, VisRet with top-1 retrieval outperforms the original-query setting even with top-10 retrieval, highlighting its ability to retrieve images that contain the required features with high precision.
+Overall, these results show that improvements in retrieval quality translate into tangible gains in downstream VQA performance.
 6 Analyses
-In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from more perspectives, including the choices of T2I instruction LLM, T2I generation model, and the downstream VQA reader.
-Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
-6.1 Query Generation Model Choice
-Does VisRet work well with other LLMs as the T2I instruction generator?
-In Table˜4, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (Grattafiori et al., 2024).
-Overall, we observe promising results: For all three LLMs, using the LLMs themselves to generate T2I instructions for VisRet outperforms using the LLM themselves to rephrase the query.
-While more expensive LLMs achieve higher performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
+6.1 Robustness across modules
+When VisRet is used in a RAG pipeline, four modules affect end-to-end performance: (1) visualization instruction generation, (2) query visualization, (3) retrieval, and (4) answer generation with an LVLM.
+We conduct analyses to probe the robustness of VisRet under variations of these modules.
+Overall, we find that changing the model for (1), (3), or (4) does not substantially alter VisRet’s effectiveness, while performance is more sensitive to (2), which depends strongly on the capability of the T2I generation model.
+We first summarize the findings below:
+  1. 1.
+    Using both 8B and 70B Llama (grattafiori2024llama) for instruction generation instead of GPT-4o, VisRet still outperforms using the LLM itself to rephrase the query (Section˜C.1).
+  2. 2.
+    While most T2I models improve over cross-modality retrieval, strong proprietary T2I generation models yield substantially larger gains when used in VisRet (Table˜3, Section˜C.2).
+We provide further discussions below.
+  3. 3.
+    VisRet continues to deliver consistent gains when using a state-of-the-art embedding model (Table˜9, Section˜C.5).
+  4. 4.
+    Repeating the VQA experiments on Visual-RAG and Visual-RAG-ME with a stronger and a weaker reader LVLM, we find that VisRet’s retrieval improvements consistently translate into downstream QA gains (Section˜C.3).
+  5. 5.
+    Although the visualized queries improve retrieval, our preliminary study suggests they still cannot replace retrieved natural images in most cases (Section˜C.4).
 Method N@1 N@10 N@30 Avg
@@
 Table 3: nDCG@k retrieval results on Visual-RAG-ME for different T2I generation models used in VisRet.
-We report N@1, N@10, N@30, and their average (Avg) for two retrievers (CLIP and E5-V).
+We report N@1, N@10, N@30, and their average.
 Best numbers within each retriever are boldfaced.
 Gemini stands for the Gemini-2.5-Flash-Image-Preview model.
+The detailed experiment setup is presented in Section˜C.2.
+T2I Model Dependency
+Despite the retrieval performance can be improved with a range of T2I generation models, VisRet is limited by the fact that it generally benefits more from stronger T2I generation models.
+As shown in Table˜3, the latest models from Google and OpenAI dominates the performance on Visual-RAG-ME by a large margin.
+Setting a lower quality for OpenAI Image-1 slightly harms the performance as well.
+What leads to the performance gap?
+In Figure˜3, we present three representative cases where Image-1 outperforms Dall-E 3 (openai_dalle3) on Visual-RAG-ME.
+We find three recurring failure modes:
+  • Lack of focus As shown in the top example, Dall-E 3 fails to zoom in and highlight the key feature wanted, i.e., a flower cluster.
+  • Factuality issues As shown in the middle example, Dall-E 3 fails to confidently render the bulb of the organism queries.
+  • Weak instruction following As shown in the bottom example, Dall-E 3 fails to comply with the instruction to show the abdomen of the organism, especially in the first visualization.
+Among them, factuality reflects a core capability gap, while the other two issues can often be partially mitigated through prompting.
+For fairness, we use the same prompt across models in Table˜3, but we expect that practical deployments can further benefit from model-specific prompt tuning.
+Figure 3: Examples of T2I generations from Image-1 and Dall-E 3 on Visual-RAG-ME.
+For these examples, VisRet with Image-1 achieves a nearly perfect retrieval score while VisRet with Dall-E 3 scores close to zero.
+[…]
+6.2 Effectiveness of Repeated Runs
+Is multi-image aggregation necessary for VisRet?
+As shown in Figure˜4, for most examples, VisRet’s N@30 outperforms at least two of its individual samples, highlighting the benefit of diversifying the query through multiple visualizations.
+At the same time, Table˜4 shows that using only one sample only slightly degrades overall performance.
+This is because in many cases, such as example 2 and 3 in Figure˜3, independently generated images share substantial information overlap.
+In such cases, a single visualization can be an efficient option.
+Figure 4: Distribution of counts where three-image VisRet outperforming its individual samples on N@30.
+Benchmark # Img N@1 N@10 N@30 Avg
+Retriever = CLIP
+Visual-RAG 3 0.251 0.431 0.438 0.373
+1 0.243 0.416 0.425 0.361
+Visual-RAG-ME 3 0.460 0.632 0.605 0.566
+1 0.449 0.623 0.602 0.558
+INQUIRE-Rerank-Hard 3 0.170 0.452 0.455 0.359
+1 0.166 0.432 0.457 0.352
+COCO-Hard 3 0.034 0.072 0.108 0.071
+1 0.033 0.068 0.105 0.069
+Retriever = E5-V
+Visual-RAG 3 0.299 0.452 0.461 0.404
+1 0.292 0.438 0.451 0.394
+Visual-RAG-ME 3 0.560 0.643 0.622 0.608
+1 0.525 0.638 0.617 0.593
+INQUIRE-Rerank-Hard 3 0.220 0.377 0.425 0.341
+1 0.205 0.292 0.426 0.308
+COCO-Hard 3 0.042 0.173 0.205 0.140
+1 0.039 0.167 0.203 0.136
+Table 4: Retrieval results (N@k) across benchmarks and image settings for two retrievers (CLIP and E5-V). Avg is the mean of N@1, N@10, and N@30.
+Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
+Visual-RAG Does the Mountain Tree Frog (scientific name: Hyla eximia) have any distinctive pattern on the underside of its body?
+[Uncaptioned image] Rank:49, nDCG@10: 0.00 [Uncaptioned image] Rank:4, nDCG@10: 0.39
+How many petals are on each of the Tower Mustard (scientific name: Turritis glabra)’s flowers? [Uncaptioned image] Rank:143, nDCG@10: 0.00 [Uncaptioned image] Rank:2, nDCG@10: 0.76
+Visual-RAG-ME Which one has striped primary flight feathers, Willet (scientific name: Tringa semipalmata) or Grey-tailed tattler (scientific name: Tringa brevipes)?
+[Uncaptioned image] Entity: Willet Rank:104, nDCG@10: 0.00 [Uncaptioned image] Entity: Willet Rank:1, nDCG@10: 0.72
+[Uncaptioned image] Entity: Grey-tailed tattler Rank:74, nDCG@10: 0.00 [Uncaptioned image] Entity: Grey-tailed tattler Rank:1, nDCG@10: 1.00
+INQUIRE A male and female cardinal sharing food [Uncaptioned image] Rank:12, nDCG@10: 0.00 [Uncaptioned image] Rank:1, nDCG@10: 1.00
+COCO Plate with pizza, knife and fork laying on edge of plate with two glasses next to them. [Uncaptioned image] Rank:39, nDCG@10: 0.00 [Uncaptioned image] Rank:3, nDCG@10: 0.3937
+Table 5: Examples: VisRet improves retrieval by highlighting visual features implied by the textual query.
+"Baseline" refers to cross-modal retrieval with the original query.
+CLIP embedding scores are used for this visualization.
+Figure 5: Examples of VisRet failing to outperform the LLM query rewriting baseline. Error patterns span T2I instruction generation, visualization, and retrieval failure.
+6.3 Success and Failure Examples
+We conclude the paper with a qualitative study.
+Table˜5 presents examples where the visualization step captures subtle visual requirements implied by the original query.
+Additional examples are provided in Appendix˜D.
+In many ground-truth images, the target entity appears in a specific posture that reveals the relevant evidence (Visual-RAG and Visual-RAG-ME), or it participates in complex relationships with other entities (INQUIRE and COCO).
+These constraints are challenging to be preserved in the embedding and surfaced during the cross-modal similarity matching.
+By contrast, VisRet incorporates them easily through query visualization, reducing the burden on the downstream embedding model.
+In Figure˜5, we show failure cases where VisRet does not outperform the LLM query rewriting baseline. We observe three salient error patterns:
+  • Instruction Generation As shown in the top row, retrieving a natural image that contains the answer requires two features: (1) the larva form and (2) the relevant body part being shown.
+The generated T2I instruction omits (1) and includes only (2).
+Addressing this issue likely requires prompt adjustments.
+  • Visualization As shown in the second row, the instruction correctly identifies that the organism’s underside must be shown, but the T2I model fails to generate a perspective that reliably highlights the feature.
+  • Retrieval Even when instruction generation and visualization succeed, retrieval can still fail when the embedding model is sensitive to variations such as background or object appearance.
+In the bottom-row example, both the background (table versus human hand and ground) and the object form (alive versus dead) differ, which leads to retrieval failure.
+7 Conclusion
+This work introduces VisRet, a novel framework that visualizes text queries to reframe cross-modality T2I retrieval as within-modality retrieval.
+By operating entirely in the visual domain, VisRet addresses key limitations of cross-modal embedding alignment.
+Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across four benchmark datasets and two retrievers.
+The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multimodal systems.
+Ethics Statement
+In this section, we describe the ethical considerations related to this paper.
+Potential Risks
+Although the goal of this paper is to introduce techniques to improve the text-to-image retrieval performance, the new approach could create new social risks.
+Specifically, in addition to the neural embedding model, our approach involves two neural models: an LLM and a T2I generation model.
+It is possible for these large models to bring in new social bias in generating the visualize query and thus bias the retrieval results.
+For instance, when depicting certain scenes of social activity, the models could reinforce stereotypical social roles.
+We urge practitioners to implement model debiasing and bias detection measure when deploying our proposed T2I retrieval method in real-world applications.
+Artifact Release
+Our Visual-RAG-ME annotation is based on Visual-RAG, which is under CC BY-NC 4.0 license and the images shared by the iNaturalist 2021 dataset, which are under one of CC BY 4.0, CC BYNC 4.0, CC BY-NC-ND 4.0, CC BY-NC-SA 4.0, CC0 1.0, CC BY-ND 4.0, CC BY-SA 4.0.
+We adhere to the intended non-commercial research use of iNaturalist 2021 dataset and do not re-distribute the images.
+Following Visual-RAG, we will release our Visual-RAG-ME annotations under CC BY-NC 4.0 license.
+Human Annotation
+Two authors, who are graduate students studying Natural Language Processing, are the only annotators involved in Visual-RAG-ME annotation.
+Both annotators are supported by the research stipend and the annotation work counted into the working hours.
+Consent was obtained from both annotators before benchmark curation.
+The entire benchmark creation process was automatically determined exempt by the institution’s IRB policy.
+The annotators actively discussed whenever they encounter ambiguity during annotation and reached agreements before proceeding.
+After the benchmark annotation, we performed a round of human auditing to ensure no question may cause privacy or ethics concerns.
+AI Assistant Use
+AI assistants, specifically ChatGPT, are used only for revising the paper draft, fixing grammar mistakes, and improving the outlook of the figures.
+Acknowledgment
+The research is supported in part by Taboola. We also thank Jia-Chen Gu and Hongwei Wang for their valuable feedback.
+Supplementary Material: Appendices
+Appendix A Benchmark Data Details
+In this section, we present the details of data processing for Visual-RAG, INQUIRE-Rerank-Hard, and COCO-Hard.
+A.1 Visual-RAG
+Visual-RAG (visual-rag) releases 391 queries with associated image names from iNaturalist 2021 (DBLP:conf/cvpr/HornCBWBA21) and corresponding retrieval labels [^2] https://github.com/visual-rag/visual-rag.
+To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test sets.
+We were able to identify all annotated images in Visual-RAG except for a single image due to a likely path error.
+A.2 INQUIRE-Rerank-Hard
+To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (DBLP:conf/nips/VendrowPSBJABH24) dataset [^3] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank.
+The original test set contained 160 queries, each paired with 100 images retrieved by CLIP.
+In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models.
+Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1.
+After manually inspecting the data, we found that for many instances, the negative images are not challenging enough and it is often very straightforward to identify the ground-truth images.
+To highlight the challenging questions, we therefore filtered out the questions for which either CLIP or E5-V achieves a perfect Recall@1.
+Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organisms themselves, with more challenging confounding negative images.
+A.3 COCO-Hard
+We derive COCO-Hard from the publicly released validation set of Microsoft COCO 2017 (mscoco).
+The original validation set contained 5000 queries, and we directly use the ground-truth images to construct the image corpus.
+In a pilot study, we found that CLIP is able to achieve 0.424 Recall@1 and 0.756 Recall@10, while E5-V has 0.594 Recall@1 and 0.892 Recall@10.
+A manual analysis shows that the queries of most examples in the original COCO dataset are short phrases centered on a single entity.
+To increase the task’s difficulty, we rank the examples in reverse order by nDCG@30 and retain the hardest 500 instances.
+We find that these hard queries often describe complex spatial relationships among multiple entities.
+Appendix B Implementation Details
+In this section, we present the implementation details of VisRet and baselines.
+VisRet: T2I Generation
+To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in implies.
+The prompts for three benchmarking datasets are shown in Figure˜6, Figure˜7, and Figure˜8, respectively.
+Then, we wrap the rephrased query in a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation.
+We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high".
+For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a single instruction already results in images with a high level of diversity.
+Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
+VisRet: Retrieval
+After obtaining the generated visualizations, we encode both the visualized images and the image corpus via an off-the-shelf CLIP [^4] https://huggingface.co/openai/clip-vit-large-patch14-336 or E5-V [^5] https://huggingface.co/royokong/e5-v encoder and perform a similarity search.
+Cosine similarity is used as the similarity metric.
+For RRF, we use $\lambda=1$ to merge the rankings from multiple queries.
+All the retrieval experiments were performed on a local server with NVIDIA A100 GPU.
+VQA Answer Generation
+We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (wei2022chain).
+Concretely, the model is asked to always extract visual information, perform reasoning, and conclude their reasoning with self-verification.
+We provide the full prompt in Figure˜9 and Figure˜10.
+VQA Evaluation
+For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (visual-rag), as shown in Figure˜11.
+Since this prompt is already human-engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering.
+We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
+Baselines
+For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important features that the query is seeking.
+At an early stage of the project, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction.
+Therefore, we report the results using the same prompt for the final version for the paper.
+For the Corpus Captioning baseline, we use the Salesforce/blip-image-captioning-large model downloaded from huggingface.
+We follow the recommended parameters to run inference on local GPUs, except setting the minimum output length to 10 tokens to encourage a more detailed caption.
+For the VISA Reranking baseline, we followed the original paper’s public implementation [^6] https://github.com/XLearning-SCU/2025-ICML-VISA, using the sample prompt except changing the model to Qwen2.5-VL-72B-Instruct which we empirically find performs better.
+For the Captioning Reranking baseline, we simplify the VISA pipeline by directly prompting for a caption for each of the top retrieved images using the following prompt:
+"Write a caption that describes the appearance and posture of the main subject.
+The caption should be concise but informative so that it can be used to retrieve the images of the same subject in different postures, states, or environments."
+Retrieval Strategy LLM R@1 R@10 R@30 N@1 N@10 N@30
+Original Query - 0.210 0.583 0.737 0.210 0.355 0.385
+LLM Rewriting Llama 3.1 8B Instruct 0.238 0.563 0.737 0.238 0.365 0.385
+Llama 3.3 70B Instruct 0.240 0.575 0.742 0.240 0.377 0.399
+GPT-4o 0.238 0.586 0.737 0.238 0.360 0.395
+Visualize-then-Retrieve Llama 3.1 8B Instruct 0.243 0.606 0.780 0.243 0.405 0.428
+Llama 3.3 70B Instruct 0.256 0.627 0.790 0.256 0.413 0.437
+GPT-4o 0.251 0.645 0.793 0.251 0.431 0.438
+Table 6: Retrieval performance on Visual-RAG of with CLIP retriever, using different LLMs as T2I instruction generator for VisRet. R = Recall. N = nDCG. The best results are boldfaced.
 Knowledge Visual-RAG Visual-RAG-ME
@@
 Table 8: VQA performance comparison using different knowledge contexts on Visual-RAG and Visual-RAG-ME. CLIP is used as the retriever. Boldfaced numbers indicate the best in each column.
-6.2 T2I Generation Model Choice
+Appendix C Further Analyses
+In this section, we provide more comprehensive analyses to investigate the effectiveness of VisRet from additional perspectives, including the choices of T2I instruction LLM, T2I generation model, and the downstream VQA reader LVLM.
+Finally, inspired by the generative retrieval literature, we conduct a pilot study of whether the generated images could be directly used as the knowledge context for downstream question answering.
+C.1 Query Generation Model Choice
+Does VisRet work well with other LLMs as the T2I instruction generator?
+In Table˜6, we study two differently sized open-weight LLMs for rephrasing the query and generating the T2I instruction: Llama 3.1 8B Instruct and Llama 3.3 70B Instruct (grattafiori2024llama).
+Overall, we observe promising results: For all three LLMs, using the LLMs themselves to generate T2I instructions for VisRet outperforms using the LLMs themselves to rephrase the query.
+While more expensive LLMs achieve higher performance, the small 8B Llama model can already achieve decent performance at a similar level as GPT-4o.
+Retrieval Method Visual-RAG Visual-RAG-ME
+N@1 N@10 N@30 N@1 N@10 N@30
+Original Query 0.279 0.416 0.420 0.330 0.546 0.543
+LLM Rewriting 0.302 0.414 0.420 0.390 0.529 0.560
+Corpus Captioning (BLIP) 0.128 0.260 0.310 0.220 0.375 0.418
+VISA Reranking (top-30) 0.288 0.431 0.442 0.380 0.547 0.541
+Captioning Reranking (top-30) 0.280 0.465 0.496 0.140 0.250 0.257
+VisRet 0.335 0.476 0.480 0.550 0.667 0.642
+Table 9: Additional evaluation results for Retriever = RzenEmbed-v2-7B.
+nDCG (N@k) is reported.
+VisRet still strongly outperforms most baselines in most settings except captioning reranking for Visual-RAG’s N@30.
+C.2 T2I Generation Model Choice
 How strong does the T2I generation model need to be for VisRet to work well?
-We compare the default T2I generation model (Image-1 with high quality setting) with several other commonly used strong T2I generation models:
-  • Stable Diffusion 3 (Esser et al., 2024).
-  • FLUX-1.dev (Black Forest Labs, 2024).
-  • Qwen-Image (Wu et al., 2025a).
-  • DALL-E 3 API (OpenAI, 2023a).
-  • Gemini-2.5-Flash-Image-Preview API (Fortin et al., 2025).
+We compare the default T2I generation model (Image-1 with high quality setting) with several other commonly used, strong T2I generation models:
+  • Stable Diffusion 3 (DBLP:conf/icml/EsserKBEMSLLSBP24).
+  • FLUX-1.dev (bfl-announcing-2024).
+  • Qwen-Image (wu2025qwenimagetechnicalreport).
+  • DALL-E 3 API (openai_dalle3).
+  • Gemini-2.5-Flash-Image-Preview API (fortin2025-gemini25-flash-image).
   • Image-1 API with the low quality setting.
-Table˜5 shows the results with GPT-4o as the T2I instruction generation model and CLIP as the retriever model.
-We observe that while all the models can improve over cross-modality retrieval to some extent in the VisRet pipeline, their effectiveness increases as the image quality retrieves.
+Table˜3 shows the results with GPT-4o as the T2I instruction generation model and CLIP as the retriever model.
+We observe that while all the models can improve over cross-modal retrieval to some extent in the VisRet pipeline, their effectiveness increases with image quality.
 The best performance is achieved by the latest proprietary models Image-1 and Gemini-2.5-Flash-Image-Preview.
-Together, these results suggest that a good T2I generation model with strong instruction following ability is necessary for VisRet.
+Together, these results suggest that a good T2I generation model with strong instruction-following ability is necessary for VisRet.
 As further T2I generation methods improve, we anticipate that building more cost-efficient versions of VisRet is viable and promising.
-6.3 Downstream VQA Model Choice
-While we have shown the benefit of VisRet on VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities?
-In Table˜6, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14).
+C.3 Downstream VQA Model Choice
+While we have shown the benefit of VisRet for VQA for GPT-4o, does the improvement hold across LVLMs with different capabilities?
+In Table˜7, we repeat the VQA experiments with two additional LVLMs: GPT-4o-mini (version gpt-4o-mini-2024-07-18) and GPT-4.1 (version gpt-4.1-2025-04-14).
 Overall, we observe similar trends as those presented in Figure˜2.
-Both direct T2I retrieval and VisRet outperform only relying on the model’s knowledge, with VisRet substantially outperforming the former.
+Both direct T2I retrieval and VisRet outperform relying only on the model’s knowledge, with VisRet substantially outperforming the former.
 These results form the foundation for VisRet as a general plug-and-play method to enhance RAG pipelines that rely on accurate T2I retrieval.
-6.4 Image Queries as Knowledge
+C.4 Image Queries as Knowledge
 As demonstrated by previous results, a T2I generation model with strong ability to follow instructions and generate realistic images is crucial to the success of VisRet.
-A natural question is whether it is still necessary to perform retrieval instead of directly using the generated images as the knowledge?
-In Table˜7, we compare the performance of using a single image as the context with VisRet.
+A natural question is whether it is still necessary to perform retrieval rather than directly using the generated images as the knowledge?
+In Table˜8, we compare the performance of using a single image as the context with VisRet.
 Overall, we observe mixed results.
 For Visual-RAG, GPT-4o-mini achieves slightly higher performance with the generated image than top-1 retrieval, but GPT-4o and GPT-4.1 exhibit the reverse pattern.
-For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefer the generated image over top-1 retrieval (and even top-10 retrieval).
-However, when provided with top-10 retrieved images, the models generally exhibit a higher VQA performance than when using the generated image.
+For Visual-RAG-ME, both GPT-4o-mini and GPT-4.1 prefer the generated image to top-1 retrieval (and even top-10 retrieval).
+However, when provided with top-10 retrieved images, the models generally exhibit higher VQA performance than when using the generated image.
 Therefore, we conclude that retrieving natural images is still crucial for challenging VQA tasks like Visual-RAG and Visual-RAG-ME and cannot be fully replaced by pure T2I generation at this stage.
-An important further direction is to combine image generation and image retrieval to improve the quality of the retrieved knowledge.
-7 Conclusion
-This work introduces VisRet, a novel framework that visualizes text queries to reframe cross-modality T2I retrieval as within-modality retrieval.
-By operating entirely in the visual domain, VisRet addresses key limitations of cross-modal embedding alignment.
-Our experiments confirm that visualized queries substantially improve both retrieval precision and downstream VQA accuracy across four benchmark datasets and two retrievers.
-The simplicity and modularity of VisRet open up promising directions for future knowledge-intensive multimodal systems.
-Ethics Statement
-In this section, we describe the ethical considerations related to this paper.
-Potential Risks
-Although the goal of this paper is to introduce techniques to improve the text-to-image retrieval performance, the new approach could create new social risks.
-Specifically, in addition to the neural embedding model, our approach involves two neural models: an LLM and a T2I generation model.
-It is possible for these large models to bring in new social bias in generating the visualize query and thus bias the retrieval results.
-For instance, when depicting certain scenes of social activity, the models could reinforce stereotypical social roles.
-We urge practitioners to implement model debiasing and bias detection measure when deploying our proposed T2I retrieval method in real-world applications.
-Artifact Release
-Our Visual-RAG-ME annotation is based on Visual-RAG, which is under CC BY-NC 4.0 license and the images shared by the iNaturalist 2021 dataset, which are under one of CC BY 4.0, CC BYNC 4.0, CC BY-NC-ND 4.0, CC BY-NC-SA 4.0, CC0 1.0, CC BY-ND 4.0, CC BY-SA 4.0.
-We adhere to the intended non-commercial research use of iNaturalist 2021 dataset and do not re-distribute the images.
-Following Visual-RAG, we will release our Visual-RAG-ME annotations under CC BY-NC 4.0 license.
-AI Assistant Use
-AI assistants, specifically ChatGPT, are used only for revising the paper draft, fixing grammar mistakes, and improving the outlook of the figures.
-Acknowledgment
-The research is supported in part by Taboola. We also thank Jia-Chen Gu and Hongwei Wang for their valuable feedback.
-Supplementary Material: Appendices
-Appendix A Benchmark Data Details
-In this section, we present the details of data processing for Visual-RAG, INQUIRE-Rerenk-Hard, and COCO-Hard.
-A.1 Visual-RAG
-Visual-RAG (Wu et al., 2025b) releases 391 queries with associated image names from iNaturalist 2021 (Horn et al., 2021) and corresponding retrieval labels [^2] https://github.com/visual-rag/visual-rag.
-To prepare the data, we download the original iNaturalist 2021 dataset and re-collect the images from the train and test set.
-We were able to identify all annotated images in Visual-RAG except a single image due to a likely path error.
-A.2 INQUIRE-Rerank-Hard
-To prepare INQUIRE-Rerank-Hard, we accessed the publicly released INQUIRE-Rerank (Vendrow et al., 2024) benchmark [^3] https://huggingface.co/datasets/evendrow/INQUIRE-Rerank.
-The original test set contained 160 queries, each paired with 100 images retrieved by CLIP.
-In a pilot study, we tested the retrieval performance of off-the-shelf CLIP and E5-V models.
-Results showed that CLIP can achieve 0.438 Recall@1 while E5-V achieved 0.506 Recall@1.
-After manually inspecting the data, we found that for a lot of instances, the negative images are not challenging enough and it is often very straightforward to identify the ground truth images.
-To highlight the challenging questions, we therefore filtered out the questions on which either CLIP or E5-V can achieve a perfect Recall@1.
-Overall, we observe that the remaining 59 questions require more nuanced image context understanding and a higher level of knowledge of the organisms themselves, with more challenging confounder negative images.
-A.3 COCO-Hard
-We derive COCO-Hard from the publicly released validation set of Microsoft COCO 2017 (Chen et al., 2015).
-The original validation set contained 5000 queries, and we directly use the ground truth images themselves to construct the image corpus.
-In a pilot study, we found that CLIP is able to achieve 0.424 Recall@1 and 0.756 Recall@10, while E5-V has 0.594 Recall@1 and 0.892 Recall@10.
-A manual analysis shows that the queries of most of the examples in the original COCO dataset are short phrases centered on a single entity.
-To increase the task’s difficulty, we rank the examples in reverse order by nDCG@30 and retained the hardest 500 instances.
-We find these hard queries often describe complex spatial relationships among multiple entities.
-Appendix B Implementation Details
-In this section, we present the implementation details of VisRet and baselines.
-VisRet: T2I Generation
-To project the text query into the image space, we first instruct an LLM to analyze the query and highlight the key visual features in it.
-The prompts for three benchmarking datasets are shown in Figure˜3, Figure˜4, and Figure˜5, respectively.
-Then, we wrap the rephrased query in a prompt template ‘‘Generate a small image of the {rephrased_query}’’ to obtain the final instruction for T2I generation.
-We use the model gpt-4o-2024-08-06 via OpenAI API with temperature = 0 for instruction generation and the gpt-image-1 model for T2I generation with the quality flag set to "high".
-For generating multiple images for each query, we find that calling the gpt-image-1 API to return multiple images given a single instruction already results in images generated with a high level of diversity.
-Therefore, we followed this setting in this paper and save further perturbing the instruction as future work.
-VisRet: Retrieval
-After obtaining the generated visualizations, we encode both the visualized images and the image corpus via an off-the-shelf CLIP [^4] https://huggingface.co/openai/clip-vit-large-patch14-336 or E5-V [^5] https://huggingface.co/royokong/e5-v encoder and perform a similarity search.
-Cosine similarity is used as the similarity metric.
-For RRF, we used $\lambda=1$ to merge the rankings from multiple queries.
-All the retrieval experiments were performed on a local server with Nvidia A100 GPU.
-VQA Answer Generation
-We slightly modify the prompt in Visual-RAG to use chain-of-thought prompting (Wei et al., 2022).
-Concretely, the model is asked to always extract visual information, perform reasoning, and conclude its reasoning with self-verification.
-We show the detailed prompt in Figure˜6 and Figure˜7.
-VQA Evaluation
-For Visual-RAG and Visual-RAG-ME, we use the same evaluation prompt released by the authors of Visual-RAG (Wu et al., 2025b), as shown in Figure˜8.
-Since this prompt is already human engineered for evaluating more complex references and long-form answers and the answers of Visual-RAG-ME are short and easy to evaluate, we did not perform additional prompt engineering.
-We use the same prompt and gpt-4o-2024-08-06 as the LLM judge for all the VQA experiments in this paper.
-Baselines
-For the LLM rephrase baseline, we use the same prompt for VisRet T2I instruction to highlight the most important feature that the query is seeking.
-At the early stage of the project, we performed manual tuning on the prompt and found that the best-performing rephrase also serves as the best-performing T2I generation instruction.
-Therefore, we report the results with the same prompt for the final version for the paper.
-For the Corpus Captioning baseline, we use the Salesforce/blip-image-captioning-large model accessed at huggingface.
-We follow the recommended parameters to run inference on local GPUs, except setting the minimum output length to 10 tokens to encourage a more detailed caption.
-For the VISA Reranking baseline, we followed the original paper’s public implementation [^6] https://github.com/XLearning-SCU/2025-ICML-VISA, using the sample prompt except changing the model to Qwen2.5-VL-72B-Instruct which we empirically find better performance.
-For the Captioning Reranking baseline, we simplify the VISA pipeline by directly prompting for a caption for each of the top retrieved images using the following prompt:
-"Write a caption that describes the appearance and posture of the main subject.
-The caption should be concise but informative so that it can be used to retrieve the images from images of the same subject in different postures, states, or environments."
+An important direction for future work is to combine image generation and image retrieval to improve the quality of the retrieved knowledge.
+C.5 Stronger Embedding Models
+In Table˜9, we provide additional results using RzenEmbed-v2-7B [^7] https://huggingface.co/qihoo360/RzenEmbed, which is one of the best models on MMEB at the time of writing (2025 December).
+VisRet still strongly outperforms most baselines in most settings except captioning reranking for Visual-RAG’s N@30.
+This result highlights both weaker and stronger embedding models benefit from query visualization.
 Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
@@
 Table 10: Additional qualitative results on the three benchmarking datasets.
+Appendix D Further Qualitative Examples
+In Table˜10, we present additional qualitative examples on the four evaluation datasets.
+Consistent with quantitative findings, VisRet reduces the embedding model’s workload by expressing the important details in a visualization that is closer to the retrieval target, resulting in substantially better performance in terms of ground-truth rank and nDCG@10.
 You are given a query, rephrase the query into a short descriptive phrase that highlights the key part of the entity where the queried feature could be found.
@@
 Figure 11: Prompt for the LLM VQA judge used for Visual-RAG and Visual-RAG-ME.
-Appendix C Further Qualitative Examples
-In Table˜8, we present additional qualitative examples on the four evaluation datasets.
-Consistent with quantitative findings, VisRet reduces the embedding’s workload by expressing the important details in a visualization close to the retrieval target, resulting in much better performance in terms of ground truth rank and nDCG@10.